### v3

# Cohomology of modular form connections on complex curves

Abstract.
We consider the reduction cohomology of modular functions defined on complex curves via generalizations of holomorphic connections. The cohomology is explicitly found in terms of higher genus counterparts of elliptic functions as analytic continuations of solutions for functional equations. Examples of modular functions on various genera are provided.
Key words and phrases:
Cohomology; Complex curves; Modular functions; Elliptic functions; Quasi-Jacobi forms
The author state that:
1.) The paper does not contain any potential conflicts of interests.
2.) The paper does not use any datasets. No dataset were generated during and/or analysed during the current study.
3.) The paper includes all data generated or analysed during this study.
4.) Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
5.) The data of the paper can be shared openly.
6.) No AI was used to write this paper.
1. Introduction
The natural problem of computation of continuous cohomologies for non-commutative structures on manifolds has proven to be a subject of great geometrical interest [9, 23, 37, 16, 19, 47]. As it was demonstrated in [16, 47], the ordinary Gelfand-Fuks cohomology of the Lie algebra of holomorphic vector fields on complex manifolds turns to be not the most effective and general one. For Riemann surfaces, and even for higher dimension complex manifolds, the classical cohomology of vector fields becomes trivial [23]. The Lie algebra of holomorphic vector fields does not always work for cohomology. For example, it is zero for a compact Riemann surface of genus higher than one. In [16] Feigin obtained various results concerning (co)-homology of the Lie algebra cosimplicial objects of holomorphic vector fields $Lie(M)$. Taking into account results in previous approaches, it is desirable also to find a way to enrich cohomological structures. This motivates constructions of more refined cohomology description for non-commutative algebraic structures. In [9], it has been proven that the Gelfand-Fuks cohomology $H^{*}(Vect(M))$ of vector fields on a smooth compact manifold $M$ is isomorphic to the singular cohomology of the space of continuous cross sections of a certain fiber bundle over $M$.
The main aim of this paper is to introduce and compute the reduction cohomology of modular functions satisfying certain conditions related to the general form of reduction formulas [50] on complex curves [15, 6, 21, 1]. Due to the structure of modular forms [14, 7, 18] and reduction relations among them (called Zhu recursion [50] in the vertex algebra literature) used in [48, 31, 20, 45], one can form chain complexes of modular forms depending on $n$ variables that are fine enough to describe the local geometry of complex curves. The reduction formulas allow to relate a modular form depending on $n$ variables with a sum of modular forms depending on $n-1$ variables in a linear way. In the vertex algebra setting, one defines so-called $n$-point correlation functions which depend on $n$ given vertex operators [13] and $n$ ”insertion points” on a given Riemann surface and surface moduli.
In contrast to more geometrical methods, e.g., of ordinary cosimplicial cohomology for Lie algebras [16, 47], the reduction cohomology pays more attention to the analytical and modular structure of elements of chain complex spaces. Computational methods involving reduction formulas (some of them recalled in the last Section [50, 31, 20, 45]) proved their effectiveness in conformal field theory, geometrical descriptions of intertwined modules for Lie algebras, and differential geometry of integrable models.
In Section 2 we give the definition of the reduction cohomology as well as a lemma relating it to the cohomology of generalized connections on $M$. The main proposition explicitly expressing the reduction cohomology in terms of spaces of generalized elliptic functions on $M$ is proven. In Section 3 we provide motivating examples of reduction formulas for various modular functions. Results of this paper are useful for cosimplisial cohomology theory of smooth manifolds, generalizations of the Bott-Segal theorem, and have their consequences in conformal field theory [16, 47], deformation theory, non-commutative geometry, modular forms, and the theory of foliations.
2. The chain complex and cohomology
In Section 3 we collect motivating examples of meromorphic functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ subject to the general patern of the reduction formulas 2.3. The coefficients $f_{l,k,m}(z_{n+1},\mu)$ and operators $T_{l,k,m}(\mu)$ (see below), that can depend in general on an extra variable $z_{n+1}$ according to the reduction formulas, as well as a set $\mu$ of modular parameters. The explicit dependence $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ on ${\bf z}_{n}$ corresponds to a vertex operator algebra and a genus of a compact complex curve $M$. In each case described in Section 3, functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ have certain automorphic properties proven in the vertex operator algebra picture (cf. corresponding references in Section 3). For any other set of coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, one defines the corresponding chain complex and cohomology. The construction of the chain complex below has a universal manner since it does not depend on explicit modular properties of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ when they satisfy the conditions of the next Subsection. Note that the coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$ are introduced to show the general manner of reduction formulas used to construct coboundary operators, as well as to simplify notations. In some cases we can omit corresponding dependence on certain idices if they are not used in particular coefficients $f_{l,k,m}$ and operators $T_{l,k,m}$.
2.1. Chain complex spaces of $n$-variable modular forms
In this Section we introduce spaces for the chain complex of modular functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ depending on $n$ variables and satisfying certain properties on genus $g$ complex curves [11, 49, 50, 33, 34, 31, 20, 45]. Elements of spaces depend on a set of moduli parameters $\mu\in{\mathcal{M}}$ where we denote by $\mathcal{M}$ a subset of the moduli space of genus $g$ complex curve $M$. Mark $n$ points ${\bf p}_{n}=(p_{1},\ldots,p_{n})$ on a compact complex curve $M$ of genus $g$. Denote by ${\bf z}_{n}=(z_{1},\ldots,z_{n})$ local coordinates around ${\bf p}_{n}\in M$. On a complex curve $M$ of genus $g$, we consider the spaces of modular forms depending on $n$ variables
$C^{n}(\mu)=\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\},$ (2.1)
that satisfy the reduction formulas. Note that in conformal field theory, modular functions depend not only on a choice of local parameters but also on a choice of states of the theory. Consequences of such dependence are illustrated in Section 3. Given a set of coefficients $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and operators $T_{l,k,m}(\mu)$ satisfying the reduction formulas with a function $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ (see below), the co-boundary operator $\delta^{n}({\bf z}_{n+1})$ on $C^{n}(\mu)$-space is defined accordingly for $\mu$-modular functions (cf. particular examples in Appendix 3, [18, 50, 31, 20, 45]).
For $n\geq 0$, and any $z_{n+1}\in\mathbb{C}$, define $\delta^{n}:C^{n}(\mu){\rightarrow}C^{n+1}(\mu)$, given by
$\delta^{n}({\bf z}_{n+1})\;\mathcal{Z}\left({\bf z}_{n},\mu\right)=\sum\limits_{l=1}^{l(g)}\sum\limits_{k=0}^{n}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}_{n+1},\mu\right)\;T_{l,k,m}(\mu).\mathcal{Z}_{n}\left({\bf z}_{n},\mu\right),$ (2.2)
where $l(g)\geq 0$ is a constant depending on $g$, and the meaning of indices $1\leq k\leq n$, $1\leq l\leq l(g)$, $m\geq 0$ is explained below. The coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ have a specific form independent of a vertex operator algebra. For each particular genus $g\geq 0$ of $M$ and a type of modular form defined by the moduli parameter $\mu$, corresponding known operator-valued functions $f_{l,k,m}({\bf z}_{n+1},\mu)T_{l,k,m}(\mu).$ change the $k$-th argument of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ by changing $\mu$. The reduction formulas have the form:
$\mathcal{Z}\left({\bf z}_{n+1},\mu\right)=\delta^{n}({\bf z}_{n+1}).\mathcal{Z}\left({\bf z}_{n},\mu\right).$ (2.3)
For $n\geq 0$, let us denote by ${\mathfrak{Z}}_{n}$ the domain of all ${\bf z}_{n}\in\mathbb{C}^{n}$, such that the chain condition
$\delta^{n+1}({\bf z}_{n+1})\;\delta({\bf z}_{n}).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0,$ (2.4)
is satisfied for the coboundary operators (2.2) for spaces $C^{n}(\mu)$. Explicitly, the chain condition (2.4) leads to an infinite $n\geq 0$ set of equations involving functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and $\mathcal{Z}\left({\bf z}_{n},\mu\right)$:
$\sum\limits_{l^{\prime}=1\atop l=1}^{l(g)}\sum\limits_{k^{\prime}\atop k=1}^{n+1\atop n}\sum\limits_{m^{\prime}\atop m\geq 0}f_{l^{\prime},k^{\prime},m^{\prime}}\left({\bf z}_{n+2},\mu\right)f_{l,k,m}\left({\bf z}_{n+1},\mu\right)T_{l^{\prime},k^{\prime},m^{\prime}}(\mu)T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0.$ (2.5)
The spaces of modular forms satisfying the conditions (2.5) constitute a chain complex
$0\longrightarrow C^{0}\stackrel{{\scriptstyle\delta^{0}}}{{\longrightarrow}}C^{1}\stackrel{{\scriptstyle\delta^{1}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle\delta^{n-2}}}{{\longrightarrow}}C^{n-1}\stackrel{{\scriptstyle\delta^{n-1}}}{{\longrightarrow}}C^{n}\longrightarrow\ldots.$ (2.6)
For $n\geq 1$, we call the corresponding cohomology
$H^{n}(\mu)={\rm Ker}\;\delta^{n}({\bf z}_{n+1})/{\rm Im}\;\delta^{n-1}({\bf z}_{n}),$ (2.7)
the $n$-th reduction cohomology of $\mu$-modular forms on a complex curve $M$. Note that the reduction cohomology can be defined for all types of modular functions satisfying the reduction formulas (2.3). Operators $T_{l,k,m}(\mu)$, $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$, form a set of generators of an infinite-dimensional continual Lie algebra $\mathfrak{g}(\mu)$ endowed with a natural grading indexed $l$, $m$. Indeed, we set the space of functions $\mathcal{Z}({\bf z}_{n},\mu)$ as the base algebra [39, 40, 41, 46] for the continual Lie algebra $\mathfrak{g}(\mu)$, and we take its generators as $X_{k,l,m}\left(\mathcal{Z}\left({\bf z}_{n},\mu\right)\right)=T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)$, for $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$. Then the commutation relations for operators $T_{l,k,m}(\mu).$ $1\leq k\leq n$ inside $\mathcal{Z}({\bf z}_{n},\mu)$ represent the commutation relations of the continual Lie algebra $\mathfrak{g}(\mu)$. Jacobi identities for $\mathfrak{g}(\mu)$ follow from Jacobi identities of the Lie algebra of operators $T_{l,k,m}(\mu)$. For a given set of functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, and operators $T_{l,k,m}(\mu)$, the identities (2.5) resulting from equations (2.4) will be discussed in a separate paper.
2.2. The geometrical meaning of reduction formulas and conditions (2.5)
In this Section we show that the reduction formulas have the form of multipoint connections generalizing ordinary holomorphic connections on complex curves [21]. Let us define the notion of a multipoint connection which will be useful for identifying reduction cohomology in Section 2.3. Motivated by the definition of a holomorphic connection for a holomorphic bundle [21] over a smooth complex curve $M$, we introduce the definition of a multiple point connection over $M$.
Let $\mathcal{V}$ be a holomorphic vector bundle on $M$, and $M_{0}\subset M$ be its subdomain. Denote by ${\mathcal{S}\mathcal{V}}$ the space of sections of $\mathcal{V}$. A multi-point connection $\mathcal{G}$ on $\mathcal{V}$ is a $\mathbb{C}$-multi-linear map $\mathcal{G}:M^{n}\to\mathbb{C}$, such that for any holomorphic function $f$, and two sections $\phi(p)$ and $\psi(p^{\prime})$ of $\mathcal{V}$ at points $p$ and $p^{\prime}$ on $M_{0}\subset M$ correspondingly, we have
$\sum\limits_{q,q^{\prime}\in M_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)+f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right),$ (2.8)
where the summation on left hand side is performed over locuses of points $q$, $q^{\prime}$ on $M_{0}$. We denote by ${\mathcal{C}on}_{n}$ the space of connections depending on $n$ points defined over $M$.
Geometrically, for a vector bundle $\mathcal{V}$ defined over $M$, a multi-point connection (2.8) relates two sections $\phi$ and $\psi$ at points $p$ and $p^{\prime}$ with a number of sections on $M_{0}\subset M$. We call
$G(\phi,\psi)=f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)+f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)-\sum\limits_{q,q^{\prime}\in M_{0}\subset\mathcal{X}}\mathcal{G}\left(f(\psi(q^{\prime})).\phi(q)\right),$
the form of a connection $\mathcal{G}$ depending on $n$ points. The space of connection forms depending on $n$ points will be denoted by $G^{n}$. Here we prove the following
Lemma 1.
Modular functions depending on $n$ variables of the space $\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\}$ form a space of connections depending on $n$ points. For $n\geq 0$, the reduction cohomology of a compact complex curve of genus $g$ is $H^{n}(\mu)={\mathcal{C}on}^{n}/G^{n-1}$.
Proof.
For non-vanishing $f(\phi(p))$ let us write set
$\mathcal{G}=\mathcal{Z}\left({\bf z}_{n},\mu\right),\;\psi(p^{\prime})=\left({\bf z}_{n+1},\mu\right),\;\phi(p)=\left({\bf z}_{n},\mu\right),$
$\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=T^{(g)}_{k,l,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right),\;$
$f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)=-f(\phi(p))\sum\limits_{l=1}^{l(g)}f_{l,0,m}\left({\bf z}_{n+1},\mu\right)\;T_{l,0,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right),$
$\sum\limits_{q_{n},{q^{\prime}}_{n}\in\atop{\mathcal{X}}_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=f(\phi(p))\sum\limits_{k=1}^{n}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}_{n+1},\mu\right)T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right).$
Thus, the formula (2.2) gives (2.3). ∎
The geometrical meaning of (2.5) consists in the following. Due to modular properties of functions $\mathcal{Z}({\bf z}_{n},\mu)$ depending on $n$ variables, (2.5) is also interpreted as relations among modular forms. The condition (2.3) defines a complex variety in ${\bf z}_{n}\in\mathbb{C}^{n}$. As most identities (e.g., trisecant identity [12, 35] and triple product identity [31]) for modular functions depending on $n$ variables (2.5) has its algebraic-geometrical meaning. The condition (2.5) relates finite series of modular functions on $M$ with rational function coefficients (at genus $g=0$) [50], or (deformed) elliptic functions (at genus $g=1$) [50, 31], or generalizations of classical elliptic functions (at genus $g\geq 2$) [20, 45].
2.3. Cohomology
In this Section we compute the reduction cohomology defined by (2.6)–(2.7). The main result of this paper is the following
Proposition 1.
The $n$-th reduction cohomology of the spaces $C^{n}(\mu)$ (2.1) of modular forms $\mathcal{Z}({\bf z}_{n},\mu)$ is the space of recursively generated (by the reduction formulas (2.3)) functions with $z_{i}\notin{\mathfrak{Z}}_{i}$, for $1\leq i\leq n$, satisfying the condition
$\sum\limits_{l=1}^{l(g)}\sum\limits_{k=1}^{n}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}_{n+1},\mu\right)\;T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0.$ (2.9)
The first cohomology is given by the space of transversal (i.e., with vanishing sum over $q$, $q^{\prime}$) one-point connections $\mathcal{Z}\left(x_{1},\mu\right)$ provided by coefficients in terms of series of special functions. The second cohomology is given by a space of generalized higher genus complex kernels corresponding to $M$.
Proof.
By definition (2.7), the $n$-th reduction cohomology is defined by the subspace of $C^{n}(\mu)$ of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ satisfying (2.9) modulo the subspace of $C^{n}(\mu)$ modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ resulting from:
$\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ $=$ $\sum\limits_{l=1}^{l(g)}\sum\limits_{k=1}^{n-1}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}^{\prime}_{n},\mu\right)\;T_{l,k,m}(\mu).\;\mathcal{Z}\left({\bf z}^{\prime}_{n-1},\mu\right).$ (2.10)
We assume that, subject to other fixed $\mu$-parameters, modular functions depending on $n$ variables are completely determined by all choices ${\bf z}_{n}\in\mathbb{C}^{n}$. Thus, the reduction cohomology can be treated as depending on set of ${\bf z}_{n}$ only with appropriate action of endomorphisms generated by $z_{n+1}$. Consider a non-vanishing solution $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ to (2.9) for some ${\bf z}_{n}$. Let us use the reduction formulas (2.3) recursively for each $z_{i}$, $1\leq i\leq n$ of ${\bf z}_{n}$ in order to express $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ in terms of modular form without parameters $\mathcal{Z}\left(\mu\right)$, i.e., we obtain
$\mathcal{Z}\left({\bf z}_{n},\mu\right)={\mathcal{D}}({\bf z}_{n},\mu)\mathcal{Z}\left(\mu\right),$ (2.11)
as in [31]. It is clear that $z_{i}\notin{\mathfrak{Z}}_{i}$ for $1\leq i\leq n$, i.e., at each stage of the recursion procedure towards (2.11), otherwise $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ would be zero. Thus, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ is explicitly known and is repsented by a series of auxiliary functions ${\mathcal{D}}({\bf z}_{n})$ depending on moduli space parameters $\mu$. Consider now $\mathcal{Z}\left({\bf z}^{\prime}_{n}\right)$ given by (2.10). It is either vanishes when $z_{n-i}\in{\mathfrak{Z}}_{n-i}$, $2\leq i\leq n$, or given by (2.11) with ${\bf z}^{\prime}_{n}$ arguments. The general idea of deriving reduction formulas is to consider the double integration of $\mathcal{Z}\left({\bf z}_{n}\right)$ along small circles around two auxiliary variables with the action of reproduction kernels inserted. Then, these procedure leads to recursion formulas relating $\mathcal{Z}({\bf z}_{n+1},\mu)$ and $\mathcal{Z}({\bf z}_{n},\mu)$ with functional coefficients depending on the nature of corresponding modular functions, and $M$. In [48, 31] formulas to modular functions depending on $n$ variables were explicitly and recursively obtained in various specific examples. In terms of $z_{n+1}$, we are able to transfer in (2.9) the action of $T_{l,k,m}(\mu)$-operators into an analytical continuation of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ multi-valued holomorphic functions to domains $D_{n}\subset M$ with $z_{i}\neq z_{j}$ for $i\neq j$. Namely, in (2.9), the operators $T_{l,k,m}(\mu)$ shift the formal parameters ${\bf z}_{n}$ by $z_{n+1}$, i.e., ${\bf z}^{\prime}_{n}={\bf z}_{n}+z_{n+1}$. Thus, the $n$-th reduction cohomology is given by the space of analytical continuations of modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ with ${\bf z}_{n-1}\notin{\mathfrak{Z}}_{n-1}$ that are solutions of (2.9). ∎
3. Motivating examples
In this Section we provide the examples of sets of coefficients $f_{l,k,m}$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfying the reduction formulas. Functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfy certain automorphic properties with respect to corresponding groups [50, 31, 20, 45]. It is easy to see that all reduction formulas of this Section are in the form (2.2). These examples motivate the general construciton of the complex in Section 2. The reduction cohomology depends on the kind of modular forms (via moduli parameters which we denote $\mu$) and genus of $M$. As it was shown in [33, 26, 27], existence of reduction formulas is related in some sense to modularity.
3.1. The rational case
In (cf., e.g., [50]) we find for the rational case functions depending on $n$ variables, the reduction formulas
$\mathcal{Z}({\bf z}_{n+1},\mu)=\sum\limits_{k=0}^{n}\sum\limits_{m\geq 0}f_{l,k,m}(z_{n+1},z_{k})\;\mathcal{Z}({\bf z}_{n},\mu),$ (3.1)
where $f_{l,k,m}(z,w)$ is a rational function defined by
$f_{l,n,m}(z,w)=\frac{z^{-n}}{m!}\left(\frac{d}{dw}\right)^{m}\frac{w^{n}}{z-w},\;\;\iota_{z,w}f_{l,n,m}(z,w)=\sum\limits_{j\in{\mathbb{N}}}\left({n+j\atop m}\right)z^{-n-j-1}w^{n+j-1},$
where $\iota_{z,w}:\mathbb{C}[z_{1},\ldots,z_{n}]\to\mathbb{C}[[z_{1},z_{1}^{-1}\ldots,z_{n}z_{n}^{-1}]]$ are certain maps [13]. Let us take $z_{n+1}$ as the variable of expansion. Then the $n$-th reduction cohomology $H^{n}(\mu)$ is given by the space of rational functions recursively generated by (2.3) with ${\bf z}_{n}\notin{\mathfrak{Z}}_{n}$, satisfying (2.9) with rational function coefficients $f_{l,k,m}(z_{n+1},z_{k})$, and modulo the space of functions depending on $n$ variables obtained by the recursion procedure, not given by $\delta^{n-1}\mathcal{Z}({\bf z}_{n-1},\mu)$. It is possible to rewrite (2.9), in the form
$\left(\partial_{z_{n+1}}+\sum\limits_{k=1}^{n}\widetilde{f}^{(0)}_{k,m}(z_{n+1},z_{k})\right)\;\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)=0,$
which is an equation for an analytical continuation of $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)$ with different functions $\widetilde{f}_{k,m}$. Using the reduction formulas (2.3) we obtain $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)={\mathcal{D}}({\bf z}_{n+1},\mu)$, where ${\mathcal{D}}({\bf z}_{n+1},\mu)$ is given by the series of rational-valued functions in ${\bf z}_{n+1}\notin{\mathfrak{Z}}_{n}$ resulting from the recursive procedure starting from function depending on $n$ variables to the partition function. Thus, in this example, the $n$-th cohomology is the space of analytic extensions of rational function solutions to the equation (2.9) with rational function coefficients.
3.2. Modular and elliptic functions
For a variable $x$, set $D_{x}=\frac{1}{2\pi i}\partial_{x}$, and $q_{x}=e^{2\pi ix}$. Define for $m\in\mathbb{N}=\{\ell\in\mathbb{Z}:\ell>0\}$, the elliptic Weierstrass functions
$P_{1}(w,\tau)=-\sum_{n\in\mathbb{Z}\backslash\{0\}}\frac{q_{w}^{n}}{1-q^{n}}-\frac{1}{2},$
$P_{m+1}(w,\tau)=\frac{\left(-1\right)^{m}}{m!}D_{w}^{m}\left(P_{1}(w,\tau)\right)=\frac{(-1)^{m+1}}{m!}\sum_{n\in\mathbb{Z}\backslash\{0\}}\frac{n^{m}q_{w}^{n}}{1-q^{n}}.$
Next, we have the following definition. The modular Eisenstein series $E_{k}(\tau)$, defined by $E_{k}=0$ for $k$ for odd and $k\geq 2$ even $E_{k}(\tau)=-\frac{B_{k}}{k!}+\frac{2}{(k-1)!}\sum\limits_{n\geq 1}\frac{n^{k-1}q^{n}}{1-q^{n}}$, where $B_{k}$ is the $k$-th Bernoulli number defined by $(e^{z}-1)^{-1}=\displaystyle{\sum\limits_{k\geq 0}\frac{B_{k}}{k!}z^{k-1}}$. It is convenient to define $E_{0}=-1$. $E_{k}$ is a modular form for $k>2$ and a quasi-modular form for $k=2$. Therefore, $E_{k}(\gamma\tau)=(c\tau+d)^{k}E_{k}(\tau)-\delta_{k,2}\frac{c(c\tau+d)}{2\pi i}$.
For $w$, $z\in\mathbb{C}$, and $\tau\in\mathbb{H}$ let us define $\widetilde{P}_{1}(w,z,\tau)=-\sum_{n\in\mathbb{Z}}\frac{q_{w}^{n}}{1-q_{z}q^{n}}$. We also have
$\widetilde{P}_{m+1}(w,z,\tau)=\frac{(-1)^{m}}{m!}D_{w}^{m}\left(\widetilde{P}_{1}(w,z,\tau)\right)=\frac{(-1)^{m+1}}{m!}\sum_{n\in\mathbb{Z}}\frac{n^{m}q_{w}^{n}}{1-q_{z}q^{n}}.$
It is thus useful to give the following definition. For $m\in\mathbb{N}_{0}$, let
$P_{m+1,\lambda}\left(w,\tau\right)$ $=$ $\frac{(-1)^{m+1}}{m!}\sum_{n\in\mathbb{Z}\backslash\{-\lambda\}}\frac{n^{m}q_{w}^{n}}{1-q^{n+\lambda}}.$ (3.2)
One notes that $P_{1,\lambda}\left(w,\tau\right)=q_{w}^{-\lambda}(P_{1}(w,\tau)+1/2)$, with $P_{m+1,\lambda}\left(w,\tau\right)=\frac{(-1)^{m}}{m!}D_{w}^{m}\left(P_{1,\lambda}\left(w,\tau\right)\right)$. We also consider the expansion $P_{1,\lambda}(w,\tau)=\frac{1}{2\pi iw}-\sum_{k\geq 1}E_{k,\lambda}(\tau)(2\pi iw)^{k-1}$, where we find [49]
$E_{k,\lambda}(\tau)=\sum_{j=0}^{k}\frac{\lambda^{j}}{j!}E_{k-j}(\tau).$ (3.3)
We define another generating set $\widetilde{E}_{k}(z,\tau)$ for $k\geq 1$ together with $E_{2}(\tau)$ given by [36] $\widetilde{P}_{1}(w,z,\tau)=\frac{1}{2\pi iw}-\sum_{k\geq 1}\widetilde{E}_{k}(z,\tau)(2\pi iw)^{k-1}$, where we find that for $k\geq 1$,
$\widetilde{E}_{k}(z,\tau)=-\delta_{k,1}\frac{q_{z}}{q_{z}-1}-\dfrac{B_{k}}{k!}+\frac{1}{(k-1)!}\sum_{m,n\geq 1}\left(n^{k-1}q_{z}^{m}+(-1)^{k}n^{k-1}q_{z}^{-m}\right)q^{mn},$
and $\widetilde{E}_{0}(z,\tau)=-1$.
3.3. The elliptic case
Let $q=e^{2\pi i\tau}$, $q_{i}=e^{z_{i}}$, where $\tau$ is the torus modular parameter. Then the genus one Zhu recursion formula is given by the following [50]
$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)=\mathcal{Z}\left({\bf z}_{n},\mu_{0},\tau\right)+\sum\limits_{k=1}^{n}\sum\limits_{m\geq 0}P_{m+1}(z_{n+1}-z_{k},\tau)\;\mathcal{Z}({\bf z}_{n},\mu_{k,m},\tau).$ (3.4)
Here $P_{m}(z,\tau)$ denote higher Weierstrass functions defined by $P_{m}(z,\tau)=\frac{(-1)^{m}}{(m-1)!}$ $\sum\limits_{n\in{\mathbb{Z}}_{\neq 0}}\frac{n^{m-1}q_{z}^{n}}{1-q^{n}}$.
3.4. The case of deformed elliptic functions
Let $w_{n+1}\in\mathbb{R}$ and define $\phi\in U(1)$ by $\phi=\exp(2\pi i\;w_{n+1})$. For some $\theta\in U(1)$, we obtain the following generalization of Zhu’s Proposition 4.3.2 [50] for the function depending on $n$ variables [31]. Let $\theta$ and $\phi$ be as as above. Then for any ${\bf z}_{n}\in C^{n}$ we have
$\mathcal{Z}\left({\bf x}_{n+1},\mu,\tau\right)$ $=$ $\delta_{\theta,1}\delta_{\phi,1}\mathcal{Z}\left({\bf x}_{n},\mu_{0},\tau\right)$
$+$ $\sum\limits_{k=1\atop m\geq 0}^{n}p(n,k)\;P_{m+1}\left[\begin{array}[]{c}\theta\\ \phi\end{array}\right](z_{n+1}-z_{k},\tau)\;\mathcal{Z}({\bf z}_{n};\mu_{k,m},\tau),$
where $p(n,k)$ is a parity multiplier. A deformed Weierstrass function is defined as follows [10, 31]. Let $(\theta,\phi)\in U(1)\times U(1)$ denote a pair of modulus one complex parameters with $\phi=\exp(2\pi i\lambda)$ for $10\leq\lambda<$ 1. For $z\in\mathbb{C}$ and $\tau\in\mathbb{H}$ we define deformed Weierstrass functions for $k\geq 1$,
$P_{k}\left[\begin{array}[]{c}\theta\\ \phi\end{array}\right](z,\tau)=\frac{(-1)^{k}}{(k-1)!}\sum\limits_{n\in\mathbb{Z}+\lambda}^{\prime}\frac{n^{k-1}q_{z}^{n}}{1-\theta^{-1}q^{n}},$
for $q=q_{2\pi i\tau}$ where $\sum\limits^{\prime}$ means we omit $n=0$ if $(\theta,\phi)=(1,1)$.
3.5. The reduction formulas for Jacobi functions
In this Subsection we recall the reduction formulas derived in [31, 7]. For $\alpha\in\mathbb{C}$, we now provide the following reduction formula for formal Jacobi functions depending on $n$ variables. Let ${\bf z}_{n+1}\in\mathbb{C}^{n+1}$, $\alpha\in\mathbb{C}$. For $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
$\mathcal{Z}\left({\bf z}_{n+1},\mu,\tau\right)=\sum_{k=1}^{n}\sum_{m\geq 0}\widetilde{P}_{m+1}\left(\frac{z_{n+1}-z_{k}}{2\pi i},\alpha z,\tau\right)\mathcal{Z}\left({\bf z}_{n},\mu_{k,m},\tau\right).$
Recall the definition of $\widetilde{P}$. For $\alpha z=\lambda\tau+\mu\in{\mathbb{Z}\tau+\mathbb{Z}}$, we have
$\mathcal{Z}\left({\bf z}_{n+1},\mu,\tau\right)=e^{-z_{n+1}\lambda}\mathcal{Z}\left({\bf z}_{n},\mu_{0,\lambda},\tau\right)+\sum_{k=1}^{n}\sum_{m\geq 0}P_{m+1,\lambda}\left(\frac{z_{n+1}-z_{k}}{2\pi i},\tau\right)\;\mathcal{Z}\left({\bf z}_{n},\mu_{k,m},\tau\right),$
with $P_{m+1,\lambda}\left(w,\tau\right)$ defined in (3.2). Next we provide the reduction formula for Jacobi functions depending on $n$ variables. For $l\geq 1$ and $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
$\mathcal{Z}\left({\bf z}_{n+1}),\mu_{1,-l},\tau\right)=\sum_{m\geq 0}(-1)^{m+1}\binom{m+l-1}{m}\widetilde{G}_{m+l}(\alpha z,\tau)\mathcal{Z}\left({\bf z}_{n};\mu_{1,m}\tau\right)$
$\quad+\sum_{k=2}^{n}\sum_{m\geq 0}(-1)^{l+1}\binom{m+l-1}{m}\widetilde{P}_{m+l}\left(\frac{z_{1}-z_{k}}{2\pi i},\alpha z,\tau\right)\mathcal{Z}\left({\bf z}_{n},\mu_{k,m},\tau\right).$ (3.6)
The formulas above imply the next result proven in [7]. For $l\geq 1$ and $\alpha z=\lambda\tau+\mu\in{\mathbb{Z}\tau}+\mathbb{Z}$, we have
$\mathcal{Z}\left({\bf z}_{n+1},\mu_{1,-l});B\right)=(-1)^{l+1}\frac{\lambda^{l-1}}{(l-1)!}\mathcal{Z}\left({\bf z}_{n+1},\mu_{0,-1},\tau\right)$
$\quad+\sum\limits_{m\geq 0}(-1)^{m+1}\binom{m+l-1}{m}E_{m+l,\lambda}(\tau)\;\mathcal{Z}\left({\bf z}_{n},\mu_{1,m},\tau\right)$
$\quad+\sum\limits_{k=2}^{n}\sum_{m\geq 0}(-1)^{l+1}\binom{m+l-1}{m}P_{m+l,\lambda}\left(\frac{x_{1}-x_{k}}{2\pi i},\tau\right)\;\mathcal{Z}\left({\bf z}_{n},\mu_{k,m},\tau\right),$
for $E_{k,\lambda}$ given in (3.3).
3.6. Multiparameter Jacobi forms
For multiparameter Jacobi forms [11, 49, 26, 27, 7], the reduction formulas are found using an analysis that is similar to that in [50, 31]. The following two formulas reduce any multiparameter Jacobi function depending on $n$ variables to a linear combination of Jacobi functions depending on $n-1$ variables with modular coefficients. For each $1\leq j\leq m$ we have
$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)=\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\mathcal{Z}({\bf z}_{n},(\alpha)_{n},\mu(m))+\sum_{s=1}^{n}\sum_{k\geq 0}\tilde{P}_{k+1}(z_{s}-z,{\bf z}_{n}\cdot(\alpha_{n}),\tau)\mathcal{Z}({\bf z}_{n},\mu_{s,k}\tau),$
where $\delta_{{\bf z}\cdot(\mu)_{n},\mathbb{Z}}$ is $1$ if ${\bf z}_{n}\cdot(\mu)_{n}\in\mathbb{Z}$ and is 00 otherwise. With the same assumptions as above, for $p\geq 1$,
$\mathcal{Z}({\bf z}_{n+1},\mu_{1,-p},\tau)=\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\delta_{p,1}\;\mathcal{Z}({\bf z}_{n},\mu_{0},\tau)$
$\quad+(-1)^{p+1}\sum_{k\geq 0}\binom{k+p-1}{p-1}\tilde{E}_{k+p}(\tau,{\bf z}_{n}\cdot(\alpha)_{n})\mathcal{Z}({\bf z}_{n},\mu_{1,k},\tau)$
$\quad+(-1)^{p+1}\sum_{s=2}^{n}\sum_{k\geq 0}\binom{k+p-1}{p-1}\tilde{P}_{k+p}(z_{s}-z_{1},\tau,{\bf z}_{n}\cdot(\alpha)_{n})\;\mathcal{Z}({\bf z}_{n}\mu_{s,k},\tau).$
The difference of a minus sign between these equations and those found in [31] can be attributed to the minus sign difference in our definitions of the functions $P_{k}\left[\begin{smallmatrix}\zeta\\ 1\end{smallmatrix}\right](w,\tau)$ and the action of $\text{SL}_{2}(\mathbb{Z})$.
3.7. The genus two counterparts of Weierstrass functions
In this Subsection we recall the definition of genus two Weierstrass functions [20]. For $m$, $n\geq 1$, we first define a number of infinite matrices and row and column vectors: $\Gamma(m,n)=\delta_{m,-n+2p-2}$, $\Delta(m,n)=\delta_{m,n+2p-2}$. We also define the projection matrix $\Pi=\Gamma^{2}=\begin{bmatrix}\mathbbm{1}_{2p-1}&0\\ 0&\ddots\end{bmatrix}$, where ${\rm Id}_{2p-3}$ denotes the $32p-$ 3 dimensional identity matrix and ${\rm Id}_{-1}=0$. Let $\Lambda_{a}$ for $a\in\{1,2\}$ be the matrix with components
$\Lambda_{a}(m,n;\tau_{a},\epsilon)=\epsilon^{(m+n)/2}(-1)^{n+1}\binom{m+n-1}{n}E_{m+n}(\tau_{a}).$
Note that $\Lambda_{a}=SA_{a}S^{-1}$, for $A_{a}$ given by
$A_{a}=A_{a}(k,l,\tau_{a},\epsilon)=\frac{(-1)^{k+1}\epsilon^{(k+l)/2}}{\sqrt{kl}}\frac{(k+l-1)!}{(k-1)!(l-1)!}E_{k+l}(\tau_{a}).$
introduce the infinite dimensional matrices for $S$ a diagonal matrix with components $S(m,n)=\sqrt{m}\delta_{mn}$. Let $\mathbb{R}(x)$ for $x$ on the torus be the row vector with components $\mathbb{R}(x;m)=\epsilon^{\frac{m}{2}}P_{m+1}(x,\tau_{a})$, for $a\in\{1,2\}$. Let $\mathbb{X}_{a}$ be the column vector with components
$\mathbb{X}_{1}(m)=\mathbb{X}_{1}\left(m;z_{n+1},{\bf z}_{n};\mu\right)=\epsilon^{-m/2}\sum_{u\in V}\mathcal{Z}\left({\bf z}_{k},\mu_{k,m},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}_{k+1,n},\mu^{\prime},\tau_{2}\right),$
$\mathbb{X}_{2}(m)=\mathbb{X}_{2}\left(m;z_{n+1},{\bf z}_{n};\mu\right)=\epsilon^{-m/2}\sum_{u\in V}\mathcal{Z}\left({\bf x}_{k},\mu,\tau_{1}\right)\mathcal{Z}\left({\bf x}_{n-k},\mu_{n-k,m},\tau_{2}\right).$ (3.7)
Introduce also an infinite row vector $\mathbb{Q}(p;x)=\mathbb{R}(x)\Delta\left(\mathbbm{1}-\widetilde{\Lambda}_{\overline{a}}\widetilde{\Lambda}_{a}\right)^{-1}$, for $x$ on the torus. Notice that $\widetilde{\Lambda}_{a}=\Lambda_{a}\Delta$. One introduces $\mathbb{P}_{j+1}(x)$ $=$ $\frac{(-1)^{j}}{j!}$ $\mathbb{P}_{1}(x)$, and $j\geq 0$, is the column with components
$\mathbb{P}_{j+1}(x;m)=\epsilon^{\frac{m}{2}}\binom{m+j-1}{j}\left(P_{j+m}(x,\tau_{a})-\delta_{j0}E_{m}(\tau_{a})\right).$ (3.8)
One defines $\mathcal{P}_{1}(p;x,y)=\mathcal{P}_{1}(p;x,y;\tau_{1},\tau_{2},\epsilon)$, for $p\geq 1$ by
$\mathcal{P}_{1}(p;x,y)$ $=$ $P_{1}(x-y,\tau_{a})-P_{1}(x,\tau_{a})$
-- $\mathbb{Q}(p;x)\widetilde{\Lambda}_{\overline{a}}\,\mathbb{P}_{1}(y)-(1-\delta_{p1})\left(\mathbb{Q}(p;x)\Lambda_{\overline{a}}\right)(2p-2),$
for $x$, $y$ on the torus, and
$\mathcal{P}_{1}(p;x,y)$ $=$ $(-1)^{p+1}\Big{[}\mathbb{Q}(p;x)\mathbb{P}_{1}(y)+(1-\delta_{p1})\epsilon^{p-1}P_{2p-1}(x)$
$+$ $(1-\delta_{p1})\left(\mathbb{Q}(p;x)\widetilde{\Lambda}_{\overline{a}}\Lambda_{a}\right)(2p-2)\Big{]},$
for $x$ and $y$ on two torai. For $j>0$, define
$\mathcal{P}_{j+1}(p;x,y)$ $=$ $\frac{1}{j!}\partial_{y}^{j}\left(\mathcal{P}_{1}(p;x,y)\right),$
$\mathcal{P}_{j+1}(p;x,y)$ $=$ $\delta_{a,\bar{a}}P_{j+1}(x-y)+(-1)^{j+1}.\mathbb{Q}(p;x)\left(\widetilde{\Lambda}_{\overline{a}}\right)^{\delta_{a,\bar{a}}}\;\mathbb{P}_{j+1}(y).\;$ (3.9)
One calls $\mathcal{P}_{j+1}(p;x,y)$ the genus two generalized Weierstrass functions.
3.8. The genus two case
In this Subsection we recall [20] the construction and reduction formulas for modular functions defined on genus two complex curve. In particular, we use the geometric construction developed in [48]. For a complex parameter $\epsilon=z_{1}z_{2}$, the modular form with no dependence on parameters is defined on a genus two complex curve by
$\mathcal{Z}\left(\mu\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}(z_{1},\mu_{1}\tau_{1})\;\mathcal{Z}(z_{2},\mu_{2},\tau_{2}),$ (3.10)
where parameters $\mu_{1}$ and $\mu_{2}$ are related. We then recall [20] the formal genus two reduction formulas for modular functions depending on $n$ variables. Let $x_{n+1}$, ${\bf y}_{k}$ and ${\bf y}^{\prime}_{l}$ be inserted on two torai. We consider the genus two modular function depending on $n$ variables $\mathcal{Z}\left(z_{n+1},{\bf z}_{k};{\bf z}^{\prime}_{l},\mu)\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}\left(z_{n+1},{\bf x}_{k},\mu_{1},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}^{\prime}_{l},\mu_{2},\tau_{2}\right)$, where the sum as in (3.10). First, one defines the functions $\mathcal{Z}_{n,a}$ for $a\in\{1,2\}$, via elliptic quasi-modular forms
$\mathcal{Z}_{n,1}\left({\bf z}_{n+1};\mu\right)$ $=$ $\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}({\bf z}_{n+1},{\bf z}_{k}\mu_{0},\tau)\;\mathcal{Z}_{n-k}\left({\bf x}_{k+1,n},\mu^{\prime},\tau_{2}\right),$
$\mathcal{Z}_{n,2}\left({\bf z}_{n+1};\mu\right)$ $=$ $\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}_{k}\left({\bf x}_{k},\mu^{\prime},\tau_{1}\right)\;\mathcal{Z}\left(z_{n+1}),{\bf z}_{k+1,n}\right),\;\mathcal{Z}_{n,3}\left({\bf z}_{n+1};\mu\right)=\mathbb{X}_{1}^{\Pi},$
of (3.7). Let $f^{(2)}_{a}(p;z_{n+1})$, for $p\geq 1$, and $a=1$, $2$ be given by
$f^{(2)}_{a}(p;z_{n+1})=1^{\delta_{ba}}+(-1)^{p\delta_{b\overline{a}}}\epsilon^{1/2}\left(\mathbb{Q}(p;z_{n+1})\left(\widetilde{\Lambda}_{\overline{a}}\right)^{\delta_{ba}}\right)(1),$
for $z_{n+1}\in\widehat{\Sigma}^{(1)}_{b}$. Let $f^{(2)}_{3}(p;z_{n+1})$, for $z_{n+1}\in{\Sigma}^{(1)}_{a}$ be an infinite row vector given by
$f^{(2)}_{3}(p;z_{n+1})=\left(\mathbb{R}(z_{n+1})+\mathbb{Q}(p;z_{n+1})\left(\widetilde{\Lambda}_{\overline{a}}\Lambda_{a}+\Lambda_{\overline{a}}\Gamma\right)\right)\Pi.$
In [20] it is proven that the genus two function depending on $n=k+l$ variables inserted at $x_{n-k}$, ${\bf y}_{k}$ on two torai has the following reduction formula
$\mathcal{Z}({\bf x}_{n+1},\mu)=\sum\limits_{l=1}^{3}f_{l}(p;z_{n+1})\;\mathcal{Z}_{n,l}\left({\bf z}_{n+1};\mu\right)=\sum_{i=1}^{n}\sum_{j\geq 0}\mathcal{P}_{j+1}(p;z_{n+1},z_{i})\;\mathcal{Z}\left({\bf z}_{n};\mu_{i,j}\right),$
where $p$ is some parameter. with $\mathcal{P}_{j+1}(p;x,y)$ of (3.9).
3.9. The genus $g$ generalizations of elliptic functions
For purposes of the formula (3.15) we recall here certain definitions [45]. Define a column vector $X=(X_{a}(m))$, indexed by $m\geq 0$ and $a\in\mathcal{I}$ with components $X_{a}(m)=\rho_{a}^{-\frac{m}{2}}$ $\sum_{\bm{\mu_{a,m}}}$ $\mathcal{Z}(\ldots;w_{a},\mu_{a,m};\ldots)$, and a row vector $p(x)=(p_{a}(x,m))$, for $m\geq 0,a\in\mathcal{I}$ with components $p_{a}(x,m)=\rho_{a}^{\frac{m}{2}}\partial^{(0,m)}\psi_{p}^{(0)}(x,w_{a})$. Introduce the column vector $G=(G_{a}(m))$, for $m\geq 0,a\in\mathcal{I}$, given by $G=\sum_{k=1}^{n}\sum_{j\geq 0}\partial_{k}^{(j)}\;q(y_{k})\;\mathcal{Z}({\bf z}_{n},\mu_{k,j})$, where $q(y)=(q_{a}(y;m))$, for $m\geq 0$, $a\in\mathcal{I}$, is a column vector with components $q_{a}(y;m)=(-1)^{p}\rho_{a}^{\frac{m+1}{2}}\partial^{(m,0)}\psi_{p}^{(0)}(w_{-a},y)$, and $R=(R_{ab}(m,n))$, for $m$, $n\geq 0$ and $a$, $b\in\mathcal{I}$ is a doubly indexed matrix with components
$R_{ab}(m,n)=\begin{cases}(-1)^{p}\rho_{a}^{\frac{m+1}{2}}\rho_{b}^{\frac{n}{2}}\partial^{(m,n)}\psi_{p}^{(0)}(w_{-a},w_{b}),&a\neq-b,\\ (-1)^{p}\rho_{a}^{\frac{m+n+1}{2}}\mathcal{E}_{m}^{n}(w_{-a}),&a=-b,\end{cases}$ (3.11)
where $\mathcal{E}_{m}^{n}(y)=\sum_{\ell=0}^{2p-2}\partial^{(m)}f_{\ell}(y)\;\partial^{(n)}y^{\ell}$, $\psi_{p}^{(0)}(x,y)=\frac{1}{x-y}+\sum_{\ell=0}^{2p-2}f_{\ell}(x)y^{\ell}$, for any Laurent series $f_{\ell}(x)$ for $\ell=0$, $\ldots$, $22p-$ 2. Define the doubly indexed matrix $\Delta=(\Delta_{ab}(m,n))$ by $\Delta_{ab}(m,n)=\delta_{m,n+2p-1}\delta_{ab}$. Denote by $\widetilde{R}=R\Delta$, and the formal inverse $(I-\widetilde{R})^{-1}$ is given by $\left(I-\widetilde{R}\right)^{-1}=\sum_{k\geq 0}\widetilde{R}^{\,k}$. Define $\chi(x)=(\chi_{a}(x;\ell))$ and $o(\bm{y}_{k},\mu_{0})=(o_{a}(\bm{y}_{k};\mu_{0},\ell))$, are finite row and column vectors indexed by $a\in\mathcal{I}$, $20\leq\ell\leq 2p-$ 2 with $\chi_{a}(x;\ell)=\rho_{a}^{-\frac{\ell}{2}}(p(x)+\widetilde{p}(x)(I-\widetilde{R})^{-1}R)_{a}(\ell)$, $o_{a}(\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)=\rho_{a}^{\frac{\ell}{2}}X_{a}(\ell)$, and where $\widetilde{p}(x)=p(x)\Delta$. $\psi_{p}(x,y)$ is defined by $\psi_{p}(x,y)=\psi_{p}^{(0)}(x,y)+\widetilde{p}(x)(I-\widetilde{R})^{-1}q(y)$. For each $a\in\mathcal{I}_{+}$ we define a vector $\theta_{a}(x)=(\theta_{a}(x;\ell))$, indexed by $20\leq\ell\leq 2p-$ 2 with components $\theta_{a}(x;\ell)=\chi_{a}(x;\ell)+(-1)^{p}\rho_{a}^{p-1-\ell}\chi_{-a}(x;2p-2-\ell)$. Now define the following vectors of formal differential forms $P(x)=p(x)\;dx^{p}$, $Q(y)=q(y)\;dy^{1-p}$, with $\widetilde{P}(x)=P(x)\Delta$. Then with
$\Psi_{p}(x,y)=\psi_{p}(x,y)\;dx^{p}\;dy^{1-p},$ (3.12)
we have $\Psi_{p}(x,y)=\Psi_{p}^{(0)}(x,y)+\widetilde{P}(x)(I-\widetilde{R})^{-1}Q(y)$. Defining
$\Theta_{a}(x;\ell)=\theta_{a}(x;\ell)\;dx^{p},$ (3.13)
$O_{a}(\bm{y}_{k},\mu_{0},\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)\;\bm{dy}_{k}^{\beta},$ (3.14)
for some parameter $\beta$.
3.10. The genus $g$ Schottky case
In this Subsection we recall [45, 44] the construction and reduction relations for modular functions depending on $n$ variables defined on a genus $g$ Riemann surface $M$ formed in the Schottky parameterization. All expressions here are functions of formal variables $w_{\pm a}$, $\rho_{a}\in\mathbb{C}$. Then we recall the genus $g$ reduction formula with universal coefficients that have a geometrical meaning and are meromorphic on $M$. These coefficients are generalizations of the elliptic Weierstrass functions [28]. For a $2g$ local coordinates $\bm{w}_{2g}=(w_{-1},w_{1};\ldots;w_{-g},w_{g})$, of $2g$ points $(p_{-1},p_{1};\ldots;p_{-g},p_{g})$ on the Riemann sphere, consider the genus zero point function depending on $2g$ variables $\mathcal{Z}(\bm{w}_{2g},\mu)=\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)=\prod_{a\in\mathcal{I}_{+}}\rho_{a}^{\beta_{a}}\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)$, where +=$\mathcal{I}_{+}=\{1,2,\ldots,g\}$, and $\beta_{a}$ are certain parameters related to $\mu$. Let us denote $\bm{z}_{+}=(z_{1},\ldots,z_{g})$, $\bm{z}_{-}=(z_{-1},\ldots,z_{-g})$. Let $w_{a}$ for $a\in\mathcal{I}$ be $2g$ formal variables. One identify them with the canonical Schottky parameters (for details of the Schottky construction, see [45, 44]). One can define the genus $g$ modular function not depending on variables as $\mathcal{Z}=(\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{z}_{2g},\bm{w}_{2g},\mu)$, for $(\bm{w}_{2g},\bm{\rho}_{2g})=(w_{\pm 1},\rho_{1};\ldots;w_{\pm g},\rho_{g})$. Now we recall the formal reduction formulas for all genus $g$ Schottky functions depending on $n$ variables. One defines the genus $g$ formal modular function depending on $n$ variables for $\bm{y}_{n}$ by $\mathcal{Z}(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, with $\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{-1,g},\mu)$, $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{{\bf z}_{+}\in\bm{\alpha}_{g}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, where here the sum is over a basis ${\bm{\alpha}}$. It follows that $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{\bm{\alpha}_{g}\in\bm{A}}\mathcal{Z}_{\bm{\alpha}_{g}}^{(g)}(\bm{y}_{n},\mu)$, where the sum ranges over $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{g})\in\bm{A}$, for $\bm{A}=A^{\otimes{g}}$. Finally, one defines corresponding formal differential forms depending on $n$ variables $Z(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, $Z_{\bm{\alpha}_{g}}(\bm{y}_{n},\mu)=\mathcal{Z}_{\bm{\alpha}}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, where $\bm{dy^{\beta}}_{n}=\prod_{k=1}^{n}dy_{k}^{\beta_{k}}$.
In [45] is was proven that the genus $g$ formal modular differential depending on $(n+1)$ variables $Z(x;\bm{y},\mu)$, for $x_{n+1}$, and point $p_{0}$, with the coordinate $y_{n+1}$, and ${\bf p}_{n}$ with coordinates ${\bf y}_{n}$ satisfies the recursive identity for ${\bf z}_{n}=(\bm{y})$
$Z\left(x_{n+1},{\bf z}_{n},\mu\right)$ $=$ $\sum_{a=1}^{g}\Theta_{a}(y_{n+1})\;O^{W_{\bm{\alpha}}}_{a}\left(z_{n+1};{\bf z}_{n}\right)$ (3.15)
$=$ $\sum_{k=1}^{n}\sum_{j\geq 0}\partial^{(0,j)}\Psi_{p}(y_{n+1},y_{k})\mathcal{Z}\left({\bf x}_{n},\mu_{k,j}\right)dy_{k}^{j}.$
Here $\partial^{(0,j)}$ is given by $\partial^{(i,j)}f(x,y)=\partial_{x}^{(i)}\partial_{y}^{(j)}f(x,y)$, for a function $f(x,y)$, and $\partial^{(0,j)}$ denotes partial derivatives with respect to $x$ and $y_{j}$. The forms $\Psi_{p}(y_{n+1},y_{k})\;dy_{k}^{j}$ given by (3.12), where $\Theta_{a}(x)$ is of (3.13), and $O^{W_{\bm{\alpha}}}_{a}(z_{n+1},{\bf z}_{n},\mu)$ of (3.14).
Acknowledgment
The author is supported by the Institute of Mathematics, Academy of Sciences of the Czech Republic (RVO 67985840).

### v5

@@
-Cohomology of modular form connections on complex curves
+Cohomology of multipoint connections on complex curves
 Abstract.
-We consider the reduction cohomology of modular functions defined on complex curves via generalizations of holomorphic connections.
+Assuming conformal field theory correlation functions defined on complex curves satisfy the recursion formulas, we express the corresponding cohomology theory via the generalizations of holomorphic connections.
 The cohomology is explicitly found in terms of higher genus counterparts of elliptic functions as analytic continuations of solutions for functional equations.
-Examples of modular functions on various genera are provided.
+Explicit examples associated to correlation functions on various genera are provided.
 Key words and phrases:
-Cohomology; Complex curves; Modular functions; Elliptic functions; Quasi-Jacobi forms
+Cohomology; Complex curves; Correlation functions; Elliptic functions
 The author state that:
@@
 1. Introduction
-The natural problem of computation of continuous cohomologies for non-commutative structures on manifolds has proven to be a subject of great geometrical interest [9, 23, 37, 16, 19, 47].
-As it was demonstrated in [16, 47], the ordinary Gelfand-Fuks cohomology of the Lie algebra of holomorphic vector fields on complex manifolds turns to be not the most effective and general one.
-For Riemann surfaces, and even for higher dimension complex manifolds, the classical cohomology of vector fields becomes trivial [23].
-The Lie algebra of holomorphic vector fields does not always work for cohomology.
-For example, it is zero for a compact Riemann surface of genus higher than one.
-In [16] Feigin obtained various results concerning (co)-homology of the Lie algebra cosimplicial objects of holomorphic vector fields $Lie(M)$.
-Taking into account results in previous approaches, it is desirable also to find a way to enrich cohomological structures.
-This motivates constructions of more refined cohomology description for non-commutative algebraic structures.
-In [9], it has been proven that the Gelfand-Fuks cohomology $H^{*}(Vect(M))$ of vector fields on a smooth compact manifold $M$ is isomorphic to the singular cohomology of the space of continuous cross sections of a certain fiber bundle over $M$.
-The main aim of this paper is to introduce and compute the reduction cohomology of modular functions satisfying certain conditions related to the general form of reduction formulas [50] on complex curves [15, 6, 21, 1].
-Due to the structure of modular forms [14, 7, 18] and reduction relations among them (called Zhu recursion [50] in the vertex algebra literature) used in [48, 31, 20, 45], one can form chain complexes of modular forms depending on $n$ variables that are fine enough to describe the local geometry of complex curves.
-The reduction formulas allow to relate a modular form depending on $n$ variables with a sum of modular forms depending on $n-1$ variables in a linear way.
-In the vertex algebra setting, one defines so-called $n$-point correlation functions which depend on $n$ given vertex operators [13] and $n$ ”insertion points” on a given Riemann surface and surface moduli.
-In contrast to more geometrical methods, e.g., of ordinary cosimplicial cohomology for Lie algebras [16, 47], the reduction cohomology pays more attention to the analytical and modular structure of elements of chain complex spaces.
-Computational methods involving reduction formulas (some of them recalled in the last Section [50, 31, 20, 45]) proved their effectiveness in conformal field theory, geometrical descriptions of intertwined modules for Lie algebras, and differential geometry of integrable models.
-In Section 2 we give the definition of the reduction cohomology as well as a lemma relating it to the cohomology of generalized connections on $M$.
-The main proposition explicitly expressing the reduction cohomology in terms of spaces of generalized elliptic functions on $M$ is proven.
-In Section 3 we provide motivating examples of reduction formulas for various modular functions.
-Results of this paper are useful for cosimplisial cohomology theory of smooth manifolds, generalizations of the Bott-Segal theorem, and have their consequences in conformal field theory [16, 47], deformation theory, non-commutative geometry, modular forms, and the theory of foliations.
+In the literature, the question of computation of continuous cohomology of non-commutative structures associated to manifolds attracted always wide geometric interest [10, 24, 38, 17, 20, 50].
+The classical Gelfand-Fuks cohomology of the Lie algebra of holomorphic vector fields on complex manifolds does not always reaches its purposes [17, 50].
+In cases of higher dimension complex manifolds and Riemann surfaces, the classical cohomology of vector fields vanishes [24].
+(Co)-homology of the Lie algebra cosimplicial objects of holomorphic vector fields was discussed in [17].
+In [10], it has been proven that Gelfand-Fuks cohomology of vector fields on a smooth compact manifold is isomorphic to the singular cohomology of the space of continuous cross sections of a certain fiber bundle over that manifold.
+Our motivation to introduce and compute more refined cohomology description for non-commutative algebraic structures, is based on results in previous cohomology approaches.
+The main aim of this paper is to formulate and compute the recursion cohomology of correlation functions satisfying certain conditions related to the general form of recursion formulas [53] in terms of geometric multipoint connections on complex curves [16, 7, 22, 1].
+We are able to construct chain complexes of correlation functions depending on $n$ variables according with the properties of correlation functions [15, 8, 19] and corresponding Zhu recursion relations [51, 32, 21, 46, 47, 48], That properties allow us to describe the local geometry of complex curves.
+The recursion formulas allow to relate a correlation function depending on $n$ variables with a sum of correlation functions depending on n1n-1 variables in a linear way.
+In conformal field theory setting, one defines $n$-point correlation functions which depend on $n$ given vertex operators [14] and $n$ insertion points on a given Riemann surface and surface moduli.
+While more geometric methods are used in the ordinary cosimplicial cohomology for Lie algebras [17, 50], the formulation of the recursion cohomology in terms of multipoint connections introduced in this paper reflects the modular and analytic properties of the corresponding chain complex spaces.
+Computational methods based on recursion formulas (some of them recalled in the last Section [53, 32, 21, 46]) turned out very effective in conformal field theory.
+In Section ??
+we give the definition of the recursion cohomology, and provide a lemma relating it to the cohomology of generalized connections on $M$.
+The main proposition explicitly expressing the recursion cohomology in terms of spaces of generalized elliptic functions on $M$ is proven.
+In Section ??
+we provide motivating examples of recursion formulas for various correlation functions.
+A formulation of the recursion cohomology in terms of multipoint connections and Proposition ??
+have their applications in generalizations of the Bott-Segal theorem [10], computations of characteristic classes in cosimplicial cohomology theory of smooth manifolds [17, 50, 1, 7, 16, 20, 22, 24], in conformal field theory [4, 15, 18, 25, 26, 39, 44], deformation theory, non-commutative geometry, modular forms [2, 5, 6, 8, 9, 12, 13, 19, 23, 27, 28, 29, 30, 31, 34, 35, 37, 43, 52], and the theory of foliations [3].
+Taking into account the power of recursion formulas in computaions of correlation functions, Proposition ??
+allows to find explicitly the cohomology in terms of combinations of coefficient functions of the corresponding recursion formulas in particular cases mentioned in Section ??.
 2. The chain complex and cohomology
-In Section 3 we collect motivating examples of meromorphic functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ subject to the general patern of the reduction formulas 2.3.
-The coefficients $f_{l,k,m}(z_{n+1},\mu)$ and operators $T_{l,k,m}(\mu)$ (see below), that can depend in general on an extra variable $z_{n+1}$ according to the reduction formulas, as well as a set $\mu$ of modular parameters.
-The explicit dependence $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ on ${\bf z}_{n}$ corresponds to a vertex operator algebra and a genus of a compact complex curve $M$.
-In each case described in Section 3, functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ have certain automorphic properties proven in the vertex operator algebra picture (cf.
-corresponding references in Section 3).
-For any other set of coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, one defines the corresponding chain complex and cohomology.
-The construction of the chain complex below has a universal manner since it does not depend on explicit modular properties of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ when they satisfy the conditions of the next Subsection.
-Note that the coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$ are introduced to show the general manner of reduction formulas used to construct coboundary operators, as well as to simplify notations.
-In some cases we can omit corresponding dependence on certain idices if they are not used in particular coefficients $f_{l,k,m}$ and operators $T_{l,k,m}$.
-2.1. Chain complex spaces of $n$-variable modular forms
-In this Section we introduce spaces for the chain complex of modular functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ depending on $n$ variables and satisfying certain properties on genus $g$ complex curves [11, 49, 50, 33, 34, 31, 20, 45].
+2.1. Chain complex spaces of $n$-variable correlation functions
+In this Section we introduce the spaces for a chain complex of correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ depending on $n$ variables and satisfying certain properties on genus $g$ complex curves [12, 52, 53, 34, 35, 32, 21, 46, 47, 48].
 Elements of spaces depend on a set of moduli parameters $\mu\in{\mathcal{M}}$ where we denote by $\mathcal{M}$ a subset of the moduli space of genus $g$ complex curve $M$.
@@
 Denote by ${\bf z}_{n}=(z_{1},\ldots,z_{n})$ local coordinates around ${\bf p}_{n}\in M$.
-On a complex curve $M$ of genus $g$, we consider the spaces of modular forms depending on $n$ variables
-$C^{n}(\mu)=\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\},$ (2.1)
-that satisfy the reduction formulas.
-Note that in conformal field theory, modular functions depend not only on a choice of local parameters but also on a choice of states of the theory.
-Consequences of such dependence are illustrated in Section 3.
-Given a set of coefficients $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and operators $T_{l,k,m}(\mu)$ satisfying the reduction formulas with a function $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ (see below), the co-boundary operator $\delta^{n}({\bf z}_{n+1})$ on $C^{n}(\mu)$-space is defined accordingly for $\mu$-modular functions (cf.
-particular examples in Appendix 3, [18, 50, 31, 20, 45]).
+On a complex curve $M$ of genus $g$, we consider the spaces
+$C^{n}(\mu)=\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right)\right\},$ (2.1)
+of correlation functions depending on $n\geq 0$ variables.
+We assume that elements of (LABEL:cnmu) satisfy the recursion formulas.
+Note that in conformal field theory, correlation functions depend not only on a choice of local parameters but also on a choice of states of the theory.
+Consequences of such dependence are illustrated in Section ??.
+Recall that a recursion formula expresses an $n+1$-point correlation function $\mathcal{Z}\left({\bf z}_{n+1},\mu\right)$ as a linear combination of $n$-point correlation functions by means of coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and insertion operators $T_{l,k,m}(\mu)$ [53, 32, 21, 8, 46, 47, 48] Given a set of coefficients $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and operators $T_{l,k,m}(\mu)$ satisfying the recursion formulas, the co-boundary operator $\delta^{n}({\bf z}_{n+1})$ is defined on $C^{n}(\mu)$-space as an inverse map increasing the number of variables to $n$ (cf.
+explicit examples in Appendix ??, [19, 53, 32, 21, 46]).
 For $n\geq 0$, and any $z_{n+1}\in\mathbb{C}$, define $\delta^{n}:C^{n}(\mu){\rightarrow}C^{n+1}(\mu)$, given by
@@
 The coefficient functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ have a specific form independent of a vertex operator algebra.
-For each particular genus $g\geq 0$ of $M$ and a type of modular form defined by the moduli parameter $\mu$, corresponding known operator-valued functions $f_{l,k,m}({\bf z}_{n+1},\mu)T_{l,k,m}(\mu).$ change the $k$-th argument of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ by changing $\mu$.
-The reduction formulas have the form:
+For each particular genus $g\geq 0$ of $M$ and a type of correlation functions defined by the moduli parameter $\mu$, corresponding known operator-valued functions $f_{l,k,m}({\bf z}_{n+1},\mu)\;T_{l,k,m}(\mu).$ change the $k$-th argument of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ by changing $\mu$.
+The recursion formulas have the form:
 $\mathcal{Z}\left({\bf z}_{n+1},\mu\right)=\delta^{n}({\bf z}_{n+1}).\mathcal{Z}\left({\bf z}_{n},\mu\right).$ (2.3)
@@
 $\delta^{n+1}({\bf z}_{n+1})\;\delta({\bf z}_{n}).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0,$ (2.4)
-is satisfied for the coboundary operators (2.2) for spaces $C^{n}(\mu)$.
-Explicitly, the chain condition (2.4) leads to an infinite $n\geq 0$ set of equations involving functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and $\mathcal{Z}\left({\bf z}_{n},\mu\right)$:
+is satisfied for the coboundary operators (LABEL:poros) for spaces $C^{n}(\mu)$.
+The chain condition (LABEL:torba) gives an infinite $n\geq 0$ set of equations on functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$ and $\mathcal{Z}\left({\bf z}_{n},\mu\right)$:
 $\sum\limits_{l^{\prime}=1\atop l=1}^{l(g)}\sum\limits_{k^{\prime}\atop k=1}^{n+1\atop n}\sum\limits_{m^{\prime}\atop m\geq 0}f_{l^{\prime},k^{\prime},m^{\prime}}\left({\bf z}_{n+2},\mu\right)f_{l,k,m}\left({\bf z}_{n+1},\mu\right)T_{l^{\prime},k^{\prime},m^{\prime}}(\mu)T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0.$ (2.5)
-The spaces of modular forms satisfying the conditions (2.5) constitute a chain complex
-$0\longrightarrow C^{0}\stackrel{{\scriptstyle\delta^{0}}}{{\longrightarrow}}C^{1}\stackrel{{\scriptstyle\delta^{1}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle\delta^{n-2}}}{{\longrightarrow}}C^{n-1}\stackrel{{\scriptstyle\delta^{n-1}}}{{\longrightarrow}}C^{n}\longrightarrow\ldots.$ (2.6)
+The spaces of correlation functions satisfying the conditions (2.5) form the complex
+$0\longrightarrow C^{0}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{0}}}}C^{1}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{1}}}}\ldots\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{n-2}}}}C^{n-1}\mathrel{{\mathop{\longrightarrow}\limits^{\delta^{n-1}}}}C^{n}\longrightarrow\ldots.$ (2.6)
 For $n\geq 1$, we call the corresponding cohomology
 $H^{n}(\mu)={\rm Ker}\;\delta^{n}({\bf z}_{n+1})/{\rm Im}\;\delta^{n-1}({\bf z}_{n}),$ (2.7)
-the $n$-th reduction cohomology of $\mu$-modular forms on a complex curve $M$.
-Note that the reduction cohomology can be defined for all types of modular functions satisfying the reduction formulas (2.3).
+the $n$-th recursion cohomology of correlation functions on a complex curve $M$.
+Note that the recursion cohomology can be defined for all types of correlation functions satisfying the recursion formulas (LABEL:reduction).
 Operators $T_{l,k,m}(\mu)$, $0\leq l\leq l(g)$, $m\geq 0$, $1\leq k\leq n$, form a set of generators of an infinite-dimensional continual Lie algebra $\mathfrak{g}(\mu)$ endowed with a natural grading indexed $l$, $m$.
@@
 Jacobi identities for $\mathfrak{g}(\mu)$ follow from Jacobi identities of the Lie algebra of operators $T_{l,k,m}(\mu)$.
-For a given set of functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, and operators $T_{l,k,m}(\mu)$, the identities (2.5) resulting from equations (2.4) will be discussed in a separate paper.
-2.2. The geometrical meaning of reduction formulas and conditions (2.5)
-In this Section we show that the reduction formulas have the form of multipoint connections generalizing ordinary holomorphic connections on complex curves [21].
-Let us define the notion of a multipoint connection which will be useful for identifying reduction cohomology in Section 2.3.
-Motivated by the definition of a holomorphic connection for a holomorphic bundle [21] over a smooth complex curve $M$, we introduce the definition of a multiple point connection over $M$.
-Let $\mathcal{V}$ be a holomorphic vector bundle on $M$, and $M_{0}\subset M$ be its subdomain.
-Denote by ${\mathcal{S}\mathcal{V}}$ the space of sections of $\mathcal{V}$.
-A multi-point connection $\mathcal{G}$ on $\mathcal{V}$ is a $\mathbb{C}$-multi-linear map $\mathcal{G}:M^{n}\to\mathbb{C}$, such that for any holomorphic function $f$, and two sections $\phi(p)$ and $\psi(p^{\prime})$ of $\mathcal{V}$ at points $p$ and $p^{\prime}$ on $M_{0}\subset M$ correspondingly, we have
-$\sum\limits_{q,q^{\prime}\in M_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)+f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right),$ (2.8)
-where the summation on left hand side is performed over locuses of points $q$, $q^{\prime}$ on $M_{0}$.
-We denote by ${\mathcal{C}on}_{n}$ the space of connections depending on $n$ points defined over $M$.
-Geometrically, for a vector bundle $\mathcal{V}$ defined over $M$, a multi-point connection (2.8) relates two sections $\phi$ and $\psi$ at points $p$ and $p^{\prime}$ with a number of sections on $M_{0}\subset M$.
-We call
-$G(\phi,\psi)=f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)+f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)-\sum\limits_{q,q^{\prime}\in M_{0}\subset\mathcal{X}}\mathcal{G}\left(f(\psi(q^{\prime})).\phi(q)\right),$
-the form of a connection $\mathcal{G}$ depending on $n$ points. The space of connection forms depending on $n$ points will be denoted by $G^{n}$. Here we prove the following
+For a given set of functions $f_{l,k,m}\left({\bf z}_{n+1},\mu\right)$, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, and operators $T_{l,k,m}(\mu)$, the identities (2.5) resulting from equations (LABEL:torba) will be discussed in a separate paper.
+In Section ??
+we collect motivating examples of meromorphic functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ subject to the general patern of the recursion formulas (LABEL:reduction).
+The coefficients $f_{l,k,m}$ $(z_{n+1},\mu)$ and operators $T_{l,k,m}(\mu)$, that can depend in general on an extra variable $z_{n+1}$ according to the recursion formulas, as well as a set $\mu$ of modular parameters.
+The explicit dependence $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ on ${\bf z}_{n}$ corresponds to a vertex operator algebra and a genus of a compact complex curve $M$.
+In each case described in Section ??, functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ have certain automorphic properties proven in conformal field theory picture (cf.
+corresponding references in Section ??).
+For any other set of coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, one defines the corresponding chain complex and cohomology.
+The construction of the chain complex below has a universal manner since it does not depend on explicit modular properties of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ when they satisfy the conditions of the next Subsection.
+Note that the coefficients $f_{l,k,m}(z_{n+1},\mu)$, operators $T_{l,k,m}(\mu)$ are introduced to show the general manner of recursion formulas used to construct coboundary operators, as well as to simplify notations.
+In some cases we can omit corresponding dependence on certain idices if they are not used in particular coefficients $f_{l,k,m}$ and operators $T_{l,k,m}$.
+2.2. Recursion relations as mutipoint connections
+Let us now show that the recursion formulas we assume for correlation functions are represented in terms of multipoint connections.
+The notion of a multipoint connection extends usual holomorphic connections on complex curves [22].
+Let $\mathcal{E}$ and $\Omega$ bethe sheaves of holomorphic sections of a holomorphic vector bundle $\mathcal{V}\to\mathcal{X}$ and differentials defined correspondingly over a smooth complex variety $\mathcal{X}$.
+A holomorphic connection $\nabla$ on $\mathcal{E}$ is a $\mathbb{C}$-linear map $\nabla:E\to E\otimes\Omega$, which satisfy Leibniz formula $\nabla(f\phi)=\nabla f\phi+\phi\otimes dz$, for any holomorphic function $f$.
+A definition of a multipoint connection on a smooth complex curve $M$ is motivated by the definition of a holomorphic connection.
+For a holomorphic vector bundle $\mathcal{V}$ on $M$, its subdomain $M_{0}\subset M$, and the space $\mathcal{S}\mathcal{V}$} of $\mathcal{V}$-sections, introduce a multi-point connection $\mathcal{G}$ on $\mathcal{V}$ as a map $\mathcal{G}:M^{n}\to\mathbb{C}$.
+For any holomorphic function $f$, and two sections $\phi(p)$ and $\psi(p^{\prime})$ of $\mathcal{V}$ at points $p$ and $p^{\prime}$ on $M_{0}\subset M$ correspondingly, we assume that
+$f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)+f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)=\sum\limits_{q,q^{\prime}\in M_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right),$ (2.8)
+where the summation is over locuses of points $q$, $q^{\prime}$ on $M_{0}$.
+Let us denote by ${\mathcal{C}on}_{n}$ the space of multipoint connections.
+We call $G(\phi,\psi)=f(\phi(p))\;\mathcal{G}\left(\psi(p^{\prime})\right)+f(\psi(p^{\prime}))\;\mathcal{G}\left(\phi(p)\right)-\sum\limits_{q,q^{\prime}\in M_{0}\subset\mathcal{X}}\mathcal{G}\left(f(\psi(q^{\prime})).\phi(q)\right)$ the form of a connection $\mathcal{G}$.
+The space of multipoint connection forms is be denoted by $G^{n}$.
+When the sum over $q$, $q^{\prime}$ in the formulas above vanishes, we call such multipoint connections transversal.
+Next we have
 Lemma 1.
-Modular functions depending on $n$ variables of the space $\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\}$ form a space of connections depending on $n$ points.
-For $n\geq 0$, the reduction cohomology of a compact complex curve of genus $g$ is $H^{n}(\mu)={\mathcal{C}on}^{n}/G^{n-1}$.
+Correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, $n\geq 0$, belong to the space of multipoint connections.
+The corresponding recursion cohomology is given by $H^{n}(\mu)={\mathcal{C}on}^{n}/G^{n-1}$.
 Proof.
@@
 $\sum\limits_{q_{n},{q^{\prime}}_{n}\in\atop{\mathcal{X}}_{0}\subset M}\mathcal{G}\left(f(\psi(q)).\phi(q^{\prime})\right)=f(\phi(p))\sum\limits_{k=1}^{n}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}_{n+1},\mu\right)T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right).$
-Thus, the formula (2.2) gives (2.3). ∎
-The geometrical meaning of (2.5) consists in the following.
-Due to modular properties of functions $\mathcal{Z}({\bf z}_{n},\mu)$ depending on $n$ variables, (2.5) is also interpreted as relations among modular forms.
-The condition (2.3) defines a complex variety in ${\bf z}_{n}\in\mathbb{C}^{n}$.
-As most identities (e.g., trisecant identity [12, 35] and triple product identity [31]) for modular functions depending on $n$ variables (2.5) has its algebraic-geometrical meaning.
-The condition (2.5) relates finite series of modular functions on $M$ with rational function coefficients (at genus $g=0$) [50], or (deformed) elliptic functions (at genus $g=1$) [50, 31], or generalizations of classical elliptic functions (at genus $g\geq 2$) [20, 45].
-2.3. Cohomology
-In this Section we compute the reduction cohomology defined by (2.6)–(2.7). The main result of this paper is the following
+Thus, the formula (2.2) gives (LABEL:reduction). ∎
+Now let us describe the geometric meaning of the relations (2.5).
+In conformal field theory, identities coming from computations, such as the triple product identity [32] and the trisecant identity [13, 36], arrise from computations of correlation functions via recursion formulas.
+In many known examples correlation functions $\mathcal{Z}({\bf z}_{n},\mu)$ enjoy modular properties.
+Therefore, (2.5) is also interpreted as relations among modular forms.
+The condition (LABEL:reduction) defines a complex variety in ${\bf z}_{n}\in\mathbb{C}^{n}$.
+The conditions (2.5) make a connection among finite series of correlation functions on $M$ with rational function coefficients (at genus $g=0$) [53], of deformed elliptic functions (at genus $g=1$) [53, 32], or generalizations of classical elliptic functions (at genus $g\geq 2$) [21, 46].
+2.3. Computation of the recursion cohomology
+The main result of this paper is
 Proposition 1.
-The $n$-th reduction cohomology of the spaces $C^{n}(\mu)$ (2.1) of modular forms $\mathcal{Z}({\bf z}_{n},\mu)$ is the space of recursively generated (by the reduction formulas (2.3)) functions with $z_{i}\notin{\mathfrak{Z}}_{i}$, for $1\leq i\leq n$, satisfying the condition
+The $n$-th recursion cohomology of the spaces $C^{n}(\mu)$ (LABEL:cnmu) of correlation functions $\mathcal{Z}({\bf z}_{n},\mu)$ is the space of functions recursively generated by (LABEL:reduction) with $z_{i}\notin{\mathfrak{Z}}_{i}$, for $1\leq i\leq n$, and satisfying the condition
 $\sum\limits_{l=1}^{l(g)}\sum\limits_{k=1}^{n}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}_{n+1},\mu\right)\;T_{l,k,m}(\mu).\mathcal{Z}\left({\bf z}_{n},\mu\right)=0.$ (2.9)
-The first cohomology is given by the space of transversal (i.e., with vanishing sum over $q$, $q^{\prime}$) one-point connections $\mathcal{Z}\left(x_{1},\mu\right)$ provided by coefficients in terms of series of special functions.
-The second cohomology is given by a space of generalized higher genus complex kernels corresponding to $M$.
+Note that the first cohomology is given by the space of transversal one-point connections $\mathcal{Z}\left(x_{1},\mu\right)$ with coefficients in terms of special functions.
+The spaces of higher genus generalized complex reproduction kernels corresponding to $M$ represent the second cohomology.
 Proof.
-By definition (2.7), the $n$-th reduction cohomology is defined by the subspace of $C^{n}(\mu)$ of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ satisfying (2.9) modulo the subspace of $C^{n}(\mu)$ modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ resulting from:
+According definition (LABEL:pusto), the $n$-th recursion cohomology is a coset of the subspace correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ satisfying (2.9) with respect to a subspace of $C^{n}(\mu)$ of correlation functions $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ resulting from:
 $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ $=$ $\sum\limits_{l=1}^{l(g)}\sum\limits_{k=1}^{n-1}\sum\limits_{m\geq 0}f_{l,k,m}\left({\bf z}^{\prime}_{n},\mu\right)\;T_{l,k,m}(\mu).\;\mathcal{Z}\left({\bf z}^{\prime}_{n-1},\mu\right).$ (2.10)
-We assume that, subject to other fixed $\mu$-parameters, modular functions depending on $n$ variables are completely determined by all choices ${\bf z}_{n}\in\mathbb{C}^{n}$.
-Thus, the reduction cohomology can be treated as depending on set of ${\bf z}_{n}$ only with appropriate action of endomorphisms generated by $z_{n+1}$.
-Consider a non-vanishing solution $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ to (2.9) for some ${\bf z}_{n}$.
-Let us use the reduction formulas (2.3) recursively for each $z_{i}$, $1\leq i\leq n$ of ${\bf z}_{n}$ in order to express $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ in terms of modular form without parameters $\mathcal{Z}\left(\mu\right)$, i.e., we obtain
-$\mathcal{Z}\left({\bf z}_{n},\mu\right)={\mathcal{D}}({\bf z}_{n},\mu)\mathcal{Z}\left(\mu\right),$ (2.11)
-as in [31].
-It is clear that $z_{i}\notin{\mathfrak{Z}}_{i}$ for $1\leq i\leq n$, i.e., at each stage of the recursion procedure towards (2.11), otherwise $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ would be zero.
-Thus, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ is explicitly known and is repsented by a series of auxiliary functions ${\mathcal{D}}({\bf z}_{n})$ depending on moduli space parameters $\mu$.
-Consider now $\mathcal{Z}\left({\bf z}^{\prime}_{n}\right)$ given by (2.10).
-It is either vanishes when $z_{n-i}\in{\mathfrak{Z}}_{n-i}$, $2\leq i\leq n$, or given by (2.11) with ${\bf z}^{\prime}_{n}$ arguments.
-The general idea of deriving reduction formulas is to consider the double integration of $\mathcal{Z}\left({\bf z}_{n}\right)$ along small circles around two auxiliary variables with the action of reproduction kernels inserted.
-Then, these procedure leads to recursion formulas relating $\mathcal{Z}({\bf z}_{n+1},\mu)$ and $\mathcal{Z}({\bf z}_{n},\mu)$ with functional coefficients depending on the nature of corresponding modular functions, and $M$.
-In [48, 31] formulas to modular functions depending on $n$ variables were explicitly and recursively obtained in various specific examples.
-In terms of $z_{n+1}$, we are able to transfer in (2.9) the action of $T_{l,k,m}(\mu)$-operators into an analytical continuation of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ multi-valued holomorphic functions to domains $D_{n}\subset M$ with $z_{i}\neq z_{j}$ for $i\neq j$.
-Namely, in (2.9), the operators $T_{l,k,m}(\mu)$ shift the formal parameters ${\bf z}_{n}$ by $z_{n+1}$, i.e., ${\bf z}^{\prime}_{n}={\bf z}_{n}+z_{n+1}$.
-Thus, the $n$-th reduction cohomology is given by the space of analytical continuations of modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ with ${\bf z}_{n-1}\notin{\mathfrak{Z}}_{n-1}$ that are solutions of (2.9).
+With other $\mu$-parameters fixed, correlation functions are completely determined by all choices ${\bf z}_{n}\in\mathbb{C}^{n}$.
+The recursion cohomology is a function of ${\bf z}_{n}$ only with appropriate recursion formulas adition of the parameter $z_{n+1}$.
+For a non-zero solution $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ for (2.9) with some ${\bf z}_{n}$, apply the recursion formulas (LABEL:reduction) recursively for each $z_{i}$, $1\leq i\leq n$ of ${\bf z}_{n}$.
+Thus, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ is then given as an expression in the correlation function without parameters $\mathcal{Z}\left(\mu\right)$.
+We have, similar to [32],
+$\mathcal{Z}\left({\bf z}_{n},\mu\right)={\mathcal{D}}({\bf z}_{n},\mu)\mathcal{Z}\left(\mu\right).$ (2.11)
+For $z_{i}\notin{\mathfrak{Z}}_{i}$, $1\leq i\leq n$, i.e., at each step of the recursion leading to (LABEL:topaz), or, otherwise $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ vanishes.
+As a result, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ we obtain an explicitly formula a series of functions ${\mathcal{D}}({\bf z}_{n})$ of the moduli space parameters $\mu$.
+The function $\mathcal{Z}\left({\bf z}^{\prime}_{n}\right)$ given by (2.10) vanishes for $z_{n-i}\in{\mathfrak{Z}}_{n-i}$, $2\leq i\leq n$, or is commes from (LABEL:topaz) with parameters ${\bf z}^{\prime}_{n}$ .
+Recursion formulas relating the correlation functions $\mathcal{Z}({\bf z}_{n+1},\mu)$ and $\mathcal{Z}({\bf z}_{n},\mu)$ with functional coefficients were obtained as a result of the double integration of the action of reproduction kernels on $\mathcal{Z}\left({\bf z}_{n}\right)$ via small paths around two auxiliary points.
+We can move now the action of $T_{l,k,m}(\mu)$-operators in (2.9) to an analytical continuation of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ which are multivalued holomorphic functions on domains $D_{n}\subset M$ with $z_{i}\neq z_{j}$, $i\neq j$.
+In (2.9) the operators $T_{l,k,m}(\mu)$ shift the formal parameters ${\bf z}_{n}$ by $z_{n+1}$, i.e., ${\bf z}^{\prime}_{n}={\bf z}_{n}+z_{n+1}$.
+As an upshot, the space of analytical continuations of correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ with ${\bf z}_{n-1}\notin{\mathfrak{Z}}_{n-1}$ which are solutions of (2.9) gives the $n$-th recursion cohomology.
 ∎
 3. Motivating examples
-In this Section we provide the examples of sets of coefficients $f_{l,k,m}$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfying the reduction formulas.
+In this Section we provide the examples of sets of coefficients $f_{l,k,m}$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfying the recursion formulas.
 Functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfy certain automorphic properties with respect to corresponding groups [53, 32, 21, 46].
-It is easy to see that all reduction formulas of this Section are in the form (2.2).
-These examples motivate the general construciton of the complex in Section 2.
-The reduction cohomology depends on the kind of modular forms (via moduli parameters which we denote $\mu$) and genus of $M$.
-As it was shown in [33, 26, 27], existence of reduction formulas is related in some sense to modularity.
+It is easy to see that all recursion formulas of this Section are in the form (LABEL:poros).
+These examples motivate the general construciton of the complex in Section ??.
+The recursion cohomology depends on the kind of correlation functions (via moduli parameters which we denote $\mu$) and genus of $M$.
+As it was shown in [34, 27, 28], existence of recursion formulas is related in some sense to modularity.
 3.1. The rational case
-In (cf., e.g., [50]) we find for the rational case functions depending on $n$ variables, the reduction formulas
+In (cf., e.g., [53]) we find for the rational case functions depending on $n$ variables, the recursion formulas
 $\mathcal{Z}({\bf z}_{n+1},\mu)=\sum\limits_{k=0}^{n}\sum\limits_{m\geq 0}f_{l,k,m}(z_{n+1},z_{k})\;\mathcal{Z}({\bf z}_{n},\mu),$ (3.1)
@@
 Let us take $z_{n+1}$ as the variable of expansion.
-Then the $n$-th reduction cohomology $H^{n}(\mu)$ is given by the space of rational functions recursively generated by (2.3) with ${\bf z}_{n}\notin{\mathfrak{Z}}_{n}$, satisfying (2.9) with rational function coefficients $f_{l,k,m}(z_{n+1},z_{k})$, and modulo the space of functions depending on $n$ variables obtained by the recursion procedure, not given by $\delta^{n-1}\mathcal{Z}({\bf z}_{n-1},\mu)$.
+Then the $n$-th recursion cohomology $H^{n}(\mu)$ is given by the space of rational functions recursively generated by (LABEL:reduction) with ${\bf z}_{n}\notin{\mathfrak{Z}}_{n}$, satisfying (2.9) with rational function coefficients $f_{l,k,m}(z_{n+1},z_{k})$, and modulo the space of functions depending on $n$ variables obtained by the recursion procedure, not given by $\delta^{n-1}\mathcal{Z}({\bf z}_{n-1},\mu)$.
 It is possible to rewrite (2.9), in the form
@@
 which is an equation for an analytical continuation of $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)$ with different functions $\widetilde{f}_{k,m}$.
-Using the reduction formulas (2.3) we obtain $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)={\mathcal{D}}({\bf z}_{n+1},\mu)$, where ${\mathcal{D}}({\bf z}_{n+1},\mu)$ is given by the series of rational-valued functions in ${\bf z}_{n+1}\notin{\mathfrak{Z}}_{n}$ resulting from the recursive procedure starting from function depending on $n$ variables to the partition function.
+Using the recursion formulas (LABEL:reduction) we obtain $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)={\mathcal{D}}({\bf z}_{n+1},\mu)$, where ${\mathcal{D}}({\bf z}_{n+1},\mu)$ is given by the series of rational-valued functions in ${\bf z}_{n+1}\notin{\mathfrak{Z}}_{n}$ resulting from the recursive procedure starting from function depending on $n$ variables to the partition function.
 Thus, in this example, the $n$-th cohomology is the space of analytic extensions of rational function solutions to the equation (2.9) with rational function coefficients.
@@
 It is thus useful to give the following definition. For $m\in\mathbb{N}_{0}$, let
-$P_{m+1,\lambda}\left(w,\tau\right)$ $=$ $\frac{(-1)^{m+1}}{m!}\sum_{n\in\mathbb{Z}\backslash\{-\lambda\}}\frac{n^{m}q_{w}^{n}}{1-q^{n+\lambda}}.$ (3.2)
+[…]
 One notes that $P_{1,\lambda}\left(w,\tau\right)=q_{w}^{-\lambda}(P_{1}(w,\tau)+1/2)$, with $P_{m+1,\lambda}\left(w,\tau\right)=\frac{(-1)^{m}}{m!}D_{w}^{m}\left(P_{1,\lambda}\left(w,\tau\right)\right)$.
@@
 A deformed Weierstrass function is defined as follows [11, 32].
-Let $(\theta,\phi)\in U(1)\times U(1)$ denote a pair of modulus one complex parameters with $\phi=\exp(2\pi i\lambda)$ for $10\leq\lambda<$ 1.
+Let $(\theta,\phi)\in U(1)\times U(1)$ denote a pair of modulus one complex parameters with $\phi=\exp(2\pi i\lambda)$ for $0\leq\lambda<1$.
 For $z\in\mathbb{C}$ and $\tau\in\mathbb{H}$ we define deformed Weierstrass functions for $k\geq 1$,
@@
 for $q=q_{2\pi i\tau}$ where $\sum\limits^{\prime}$ means we omit $n=0$ if $(\theta,\phi)=(1,1)$.
-3.5. The reduction formulas for Jacobi functions
-In this Subsection we recall the reduction formulas derived in [31, 7].
-For $\alpha\in\mathbb{C}$, we now provide the following reduction formula for formal Jacobi functions depending on $n$ variables.
+3.5. The recursion formulas for Jacobi functions
+In this Subsection we recall the recursion formulas derived in [32, 8].
+For $\alpha\in\mathbb{C}$, we now provide the following recursion formula for formal Jacobi functions depending on $n$ variables.
 Let ${\bf z}_{n+1}\in\mathbb{C}^{n+1}$, $\alpha\in\mathbb{C}$.
@@
 with $P_{m+1,\lambda}\left(w,\tau\right)$ defined in (3.2).
-Next we provide the reduction formula for Jacobi functions depending on $n$ variables.
+Next we provide the recursion formula for Jacobi functions depending on $n$ variables.
 For $l\geq 1$ and $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
@@
 $\quad+\sum\limits_{k=2}^{n}\sum_{m\geq 0}(-1)^{l+1}\binom{m+l-1}{m}P_{m+l,\lambda}\left(\frac{x_{1}-x_{k}}{2\pi i},\tau\right)\;\mathcal{Z}\left({\bf z}_{n},\mu_{k,m},\tau\right),$
-for $E_{k,\lambda}$ given in (3.3).
+for $E_{k,\lambda}$ given in (LABEL:eq:Gkl).
 3.6. Multiparameter Jacobi forms
-For multiparameter Jacobi forms [11, 49, 26, 27, 7], the reduction formulas are found using an analysis that is similar to that in [50, 31].
-The following two formulas reduce any multiparameter Jacobi function depending on $n$ variables to a linear combination of Jacobi functions depending on $n-1$ variables with modular coefficients.
+For multiparameter Jacobi forms [12, 52, 27, 28, 8], the recursion formulas are found using an analysis that is similar to that in [53, 32].
+The following two formulas reduce any multiparameter Jacobi function depending on $n$ variables to a linear combination of Jacobi functions depending on n1n-1 variables with modular coefficients.
 For each $1\leq j\leq m$ we have
-$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)=\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\mathcal{Z}({\bf z}_{n},(\alpha)_{n},\mu(m))+\sum_{s=1}^{n}\sum_{k\geq 0}\tilde{P}_{k+1}(z_{s}-z,{\bf z}_{n}\cdot(\alpha_{n}),\tau)\mathcal{Z}({\bf z}_{n},\mu_{s,k}\tau),$
-where $\delta_{{\bf z}\cdot(\mu)_{n},\mathbb{Z}}$ is $1$ if ${\bf z}_{n}\cdot(\mu)_{n}\in\mathbb{Z}$ and is 00 otherwise. With the same assumptions as above, for $p\geq 1$,
+$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)$ $=$ $\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\mathcal{Z}({\bf z}_{n},(\alpha)_{n},\mu(m))$
+$+$ $\sum_{s=1}^{n}\sum_{k\geq 0}\tilde{P}_{k+1}(z_{s}-z,{\bf z}_{n}\cdot(\alpha_{n}),\tau)\mathcal{Z}({\bf z}_{n},\mu_{s,k}\tau),$
+where $\delta_{{\bf z}\cdot(\mu)_{n},\mathbb{Z}}$ is $1$ if ${\bf z}_{n}\cdot(\mu)_{n}\in\mathbb{Z}$ and is 0 otherwise. With the same assumptions as above, for $p\geq 1$,
 $\mathcal{Z}({\bf z}_{n+1},\mu_{1,-p},\tau)=\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\delta_{p,1}\;\mathcal{Z}({\bf z}_{n},\mu_{0},\tau)$
@@
 For $m$, $n\geq 1$, we first define a number of infinite matrices and row and column vectors: $\Gamma(m,n)=\delta_{m,-n+2p-2}$, $\Delta(m,n)=\delta_{m,n+2p-2}$.
-We also define the projection matrix $\Pi=\Gamma^{2}=\begin{bmatrix}\mathbbm{1}_{2p-1}&0\\ 0&\ddots\end{bmatrix}$, where ${\rm Id}_{2p-3}$ denotes the $32p-$ 3 dimensional identity matrix and ${\rm Id}_{-1}=0$.
+We also define the projection matrix $\Pi=\Gamma^{2}=\begin{bmatrix}\mathbbm{1}_{2p-1}&0\\ 0&\ddots\end{bmatrix}$, where ${\rm Id}_{2p-3}$ denotes the 2p32p-3 dimensional identity matrix and ${\rm Id}_{-1}=0$.
 Let $\Lambda_{a}$ for $a\in\{1,2\}$ be the matrix with components
@@
 Notice that $\widetilde{\Lambda}_{a}=\Lambda_{a}\Delta$.
-One introduces $\mathbb{P}_{j+1}(x)$ $=$ $\frac{(-1)^{j}}{j!}$ $\mathbb{P}_{1}(x)$, and $j\geq 0$, is the column with components
+One introduces $\mathbb{P}_{j+1}(x)$ == $\frac{(-1)^{j}}{j!}$ $\mathbb{P}_{1}(x)$, and $j\geq 0$, is the column with components
 $\mathbb{P}_{j+1}(x;m)=\epsilon^{\frac{m}{2}}\binom{m+j-1}{j}\left(P_{j+m}(x,\tau_{a})-\delta_{j0}E_{m}(\tau_{a})\right).$ (3.8)
@@
 3.8. The genus two case
-In this Subsection we recall [20] the construction and reduction formulas for modular functions defined on genus two complex curve.
+In this Subsection we recall [21] the construction and recursion formulas for correlation functions defined on genus two complex curve.
 In particular, we use the geometric construction developed in [51].
-For a complex parameter $\epsilon=z_{1}z_{2}$, the modular form with no dependence on parameters is defined on a genus two complex curve by
+For a complex parameter $\epsilon=z_{1}z_{2}$, the correlation function with no dependence on parameters is defined on a genus two complex curve by
 $\mathcal{Z}\left(\mu\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}(z_{1},\mu_{1}\tau_{1})\;\mathcal{Z}(z_{2},\mu_{2},\tau_{2}),$ (3.10)
 where parameters $\mu_{1}$ and $\mu_{2}$ are related.
-We then recall [20] the formal genus two reduction formulas for modular functions depending on $n$ variables.
+We then recall [21] the formal genus two recursion formulas for correlation functions depending on $n$ variables.
 Let $x_{n+1}$, ${\bf y}_{k}$ and ${\bf y}^{\prime}_{l}$ be inserted on two torai.
-We consider the genus two modular function depending on $n$ variables $\mathcal{Z}\left(z_{n+1},{\bf z}_{k};{\bf z}^{\prime}_{l},\mu)\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}\left(z_{n+1},{\bf x}_{k},\mu_{1},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}^{\prime}_{l},\mu_{2},\tau_{2}\right)$, where the sum as in (3.10).
+We consider the genus two correlation function depending on $n$ variables $\mathcal{Z}\left(z_{n+1},{\bf z}_{k};{\bf z}^{\prime}_{l},\mu)\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}\left(z_{n+1},{\bf x}_{k},\mu_{1},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}^{\prime}_{l},\mu_{2},\tau_{2}\right)$, where the sum as in (LABEL:eq:Zdef).
 First, one defines the functions $\mathcal{Z}_{n,a}$ for $a\in\{1,2\}$, via elliptic quasi-modular forms
@@
 $\mathcal{Z}_{n,2}\left({\bf z}_{n+1};\mu\right)$ $=$ $\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}_{k}\left({\bf x}_{k},\mu^{\prime},\tau_{1}\right)\;\mathcal{Z}\left(z_{n+1}),{\bf z}_{k+1,n}\right),\;\mathcal{Z}_{n,3}\left({\bf z}_{n+1};\mu\right)=\mathbb{X}_{1}^{\Pi},$
-of (3.7). Let $f^{(2)}_{a}(p;z_{n+1})$, for $p\geq 1$, and $a=1$, $2$ be given by
+of (3.7). Let $f^{(2)}_{a}(p;z_{n+1})$, for $p\geq 1$, and $a=1$, 22 be given by
 $f^{(2)}_{a}(p;z_{n+1})=1^{\delta_{ba}}+(-1)^{p\delta_{b\overline{a}}}\epsilon^{1/2}\left(\mathbb{Q}(p;z_{n+1})\left(\widetilde{\Lambda}_{\overline{a}}\right)^{\delta_{ba}}\right)(1),$
@@
 $f^{(2)}_{3}(p;z_{n+1})=\left(\mathbb{R}(z_{n+1})+\mathbb{Q}(p;z_{n+1})\left(\widetilde{\Lambda}_{\overline{a}}\Lambda_{a}+\Lambda_{\overline{a}}\Gamma\right)\right)\Pi.$
-In [20] it is proven that the genus two function depending on $n=k+l$ variables inserted at $x_{n-k}$, ${\bf y}_{k}$ on two torai has the following reduction formula
+In [21] it is proven that the genus two function depending on $n=k+l$ variables inserted at $x_{n-k}$, ${\bf y}_{k}$ on two torai has the following recursion formula
 $\mathcal{Z}({\bf x}_{n+1},\mu)=\sum\limits_{l=1}^{3}f_{l}(p;z_{n+1})\;\mathcal{Z}_{n,l}\left({\bf z}_{n+1};\mu\right)=\sum_{i=1}^{n}\sum_{j\geq 0}\mathcal{P}_{j+1}(p;z_{n+1},z_{i})\;\mathcal{Z}\left({\bf z}_{n};\mu_{i,j}\right),$
@@
 $R_{ab}(m,n)=\begin{cases}(-1)^{p}\rho_{a}^{\frac{m+1}{2}}\rho_{b}^{\frac{n}{2}}\partial^{(m,n)}\psi_{p}^{(0)}(w_{-a},w_{b}),&a\neq-b,\\ (-1)^{p}\rho_{a}^{\frac{m+n+1}{2}}\mathcal{E}_{m}^{n}(w_{-a}),&a=-b,\end{cases}$ (3.11)
-where $\mathcal{E}_{m}^{n}(y)=\sum_{\ell=0}^{2p-2}\partial^{(m)}f_{\ell}(y)\;\partial^{(n)}y^{\ell}$, $\psi_{p}^{(0)}(x,y)=\frac{1}{x-y}+\sum_{\ell=0}^{2p-2}f_{\ell}(x)y^{\ell}$, for any Laurent series $f_{\ell}(x)$ for $\ell=0$, $\ldots$, $22p-$ 2.
+where $\mathcal{E}_{m}^{n}(y)=\sum_{\ell=0}^{2p-2}\partial^{(m)}f_{\ell}(y)\;\partial^{(n)}y^{\ell}$, $\psi_{p}^{(0)}(x,y)=\frac{1}{x-y}+\sum_{\ell=0}^{2p-2}f_{\ell}(x)y^{\ell}$, for any Laurent series $f_{\ell}(x)$ for $\ell=0$, $\ldots$, 2p22p-2.
 Define the doubly indexed matrix $\Delta=(\Delta_{ab}(m,n))$ by $\Delta_{ab}(m,n)=\delta_{m,n+2p-1}\delta_{ab}$.
 Denote by $\widetilde{R}=R\Delta$, and the formal inverse $(I-\widetilde{R})^{-1}$ is given by $\left(I-\widetilde{R}\right)^{-1}=\sum_{k\geq 0}\widetilde{R}^{\,k}$.
-Define $\chi(x)=(\chi_{a}(x;\ell))$ and $o(\bm{y}_{k},\mu_{0})=(o_{a}(\bm{y}_{k};\mu_{0},\ell))$, are finite row and column vectors indexed by $a\in\mathcal{I}$, $20\leq\ell\leq 2p-$ 2 with $\chi_{a}(x;\ell)=\rho_{a}^{-\frac{\ell}{2}}(p(x)+\widetilde{p}(x)(I-\widetilde{R})^{-1}R)_{a}(\ell)$, $o_{a}(\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)=\rho_{a}^{\frac{\ell}{2}}X_{a}(\ell)$, and where $\widetilde{p}(x)=p(x)\Delta$.
+Define $\chi(x)=(\chi_{a}(x;\ell))$ and $o(\bm{y}_{k},\mu_{0})=(o_{a}(\bm{y}_{k};\mu_{0},\ell))$, are finite row and column vectors indexed by $a\in\mathcal{I}$, $0\leq\ell\leq 2p-2$ with $\chi_{a}(x;\ell)=\rho_{a}^{-\frac{\ell}{2}}(p(x)+\widetilde{p}(x)(I-\widetilde{R})^{-1}R)_{a}(\ell)$, $o_{a}(\ell)=o_{a}(\bm{y}_{k},\mu_{0},\ell)=\rho_{a}^{\frac{\ell}{2}}X_{a}(\ell)$, and where $\widetilde{p}(x)=p(x)\Delta$.
 $\psi_{p}(x,y)$ is defined by $\psi_{p}(x,y)=\psi_{p}^{(0)}(x,y)+\widetilde{p}(x)(I-\widetilde{R})^{-1}q(y)$.
-For each $a\in\mathcal{I}_{+}$ we define a vector $\theta_{a}(x)=(\theta_{a}(x;\ell))$, indexed by $20\leq\ell\leq 2p-$ 2 with components $\theta_{a}(x;\ell)=\chi_{a}(x;\ell)+(-1)^{p}\rho_{a}^{p-1-\ell}\chi_{-a}(x;2p-2-\ell)$.
+For each $a\in\mathcal{I}_{+}$ we define a vector $\theta_{a}(x)=(\theta_{a}(x;\ell))$, indexed by $0\leq\ell\leq 2p-2$ with components $\theta_{a}(x;\ell)=\chi_{a}(x;\ell)+(-1)^{p}\rho_{a}^{p-1-\ell}\chi_{-a}(x;2p-2-\ell)$.
 Now define the following vectors of formal differential forms $P(x)=p(x)\;dx^{p}$, $Q(y)=q(y)\;dy^{1-p}$, with $\widetilde{P}(x)=P(x)\Delta$.
@@
 3.10. The genus $g$ Schottky case
-In this Subsection we recall [45, 44] the construction and reduction relations for modular functions depending on $n$ variables defined on a genus $g$ Riemann surface $M$ formed in the Schottky parameterization.
+In this Subsection we recall [46, 45] the construction and recursion relations for correlation functions depending on $n$ variables defined on a genus $g$ Riemann surface $M$ formed in the Schottky parameterization.
 All expressions here are functions of formal variables $w_{\pm a}$, $\rho_{a}\in\mathbb{C}$.
-Then we recall the genus $g$ reduction formula with universal coefficients that have a geometrical meaning and are meromorphic on $M$.
+Then we recall the genus $g$ recursion formula with universal coefficients that have a geometric meaning and are meromorphic on $M$.
 These coefficients are generalizations of the elliptic Weierstrass functions [29].
-For a $2g$ local coordinates $\bm{w}_{2g}=(w_{-1},w_{1};\ldots;w_{-g},w_{g})$, of $2g$ points $(p_{-1},p_{1};\ldots;p_{-g},p_{g})$ on the Riemann sphere, consider the genus zero point function depending on $2g$ variables $\mathcal{Z}(\bm{w}_{2g},\mu)=\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)=\prod_{a\in\mathcal{I}_{+}}\rho_{a}^{\beta_{a}}\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)$, where +=$\mathcal{I}_{+}=\{1,2,\ldots,g\}$, and $\beta_{a}$ are certain parameters related to $\mu$.
+For a $2g$ local coordinates $\bm{w}_{2g}=(w_{-1},w_{1};\ldots;w_{-g},w_{g})$, of $2g$ points $(p_{-1},p_{1};\ldots;p_{-g},p_{g})$ on the Riemann sphere, consider the genus zero point function depending on $2g$ variables $\mathcal{Z}(\bm{w}_{2g},\mu)=\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)=\prod_{a\in\mathcal{I}_{+}}\rho_{a}^{\beta_{a}}\mathcal{Z}(w_{-1},w_{1};\ldots;w_{-g},w_{g},\mu)$, where $\mathcal{I}_{+}=\{1,2,\ldots,g\}$, and $\beta_{a}$ are certain parameters related to $\mu$.
 Let us denote $\bm{z}_{+}=(z_{1},\ldots,z_{g})$, $\bm{z}_{-}=(z_{-1},\ldots,z_{-g})$.
@@
 One identify them with the canonical Schottky parameters (for details of the Schottky construction, see [46, 45]).
-One can define the genus $g$ modular function not depending on variables as $\mathcal{Z}=(\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{z}_{2g},\bm{w}_{2g},\mu)$, for $(\bm{w}_{2g},\bm{\rho}_{2g})=(w_{\pm 1},\rho_{1};\ldots;w_{\pm g},\rho_{g})$.
-Now we recall the formal reduction formulas for all genus $g$ Schottky functions depending on $n$ variables.
-One defines the genus $g$ formal modular function depending on $n$ variables for $\bm{y}_{n}$ by $\mathcal{Z}(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, with $\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{-1,g},\mu)$, $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{{\bf z}_{+}\in\bm{\alpha}_{g}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, where here the sum is over a basis ${\bm{\alpha}}$.
+One can define the genus $g$ correlation function not depending on variables as $\mathcal{Z}=(\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{z}_{2g},\bm{w}_{2g},\mu)$, for $(\bm{w}_{2g},\bm{\rho}_{2g})=(w_{\pm 1},\rho_{1};\ldots;w_{\pm g},\rho_{g})$.
+Now we recall the formal recursion formulas for all genus $g$ Schottky functions depending on $n$ variables.
+One defines the genus $g$ formal correlation function depending on $n$ variables for $\bm{y}_{n}$ by $\mathcal{Z}(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\bm{\rho}_{2g},\mu)=\sum_{\bm{z}_{+}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, with $\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)=\mathcal{Z}(\bm{y}_{n};\bm{w}_{-1,g},\mu)$, $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{{\bf z}_{+}\in\bm{\alpha}_{g}}\mathcal{Z}(\bm{y}_{n};\bm{w}_{2g},\mu)$, where here the sum is over a basis $\bm{\alpha}$}.
 It follows that $\mathcal{Z}(\bm{y}_{n},\mu)=\sum_{\bm{\alpha}_{g}\in\bm{A}}\mathcal{Z}_{\bm{\alpha}_{g}}^{(g)}(\bm{y}_{n},\mu)$, where the sum ranges over $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{g})\in\bm{A}$, for $\bm{A}=A^{\otimes{g}}$.
 Finally, one defines corresponding formal differential forms depending on $n$ variables $Z(\bm{y}_{n},\mu)=\mathcal{Z}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, $Z_{\bm{\alpha}_{g}}(\bm{y}_{n},\mu)=\mathcal{Z}_{\bm{\alpha}}(\bm{y}_{n},\mu)\;\bm{dy^{\beta}}_{n}$, where $\bm{dy^{\beta}}_{n}=\prod_{k=1}^{n}dy_{k}^{\beta_{k}}$.
-In [45] is was proven that the genus $g$ formal modular differential depending on $(n+1)$ variables $Z(x;\bm{y},\mu)$, for $x_{n+1}$, and point $p_{0}$, with the coordinate $y_{n+1}$, and ${\bf p}_{n}$ with coordinates ${\bf y}_{n}$ satisfies the recursive identity for ${\bf z}_{n}=(\bm{y})$
+In [46] is was proven that the genus $g$ formal correlation differential depending on $(n+1)$ variables $Z(x;\bm{y},\mu)$, for $x_{n+1}$, and point $p_{0}$, with the coordinate $y_{n+1}$, and ${\bf p}_{n}$ with coordinates ${\bf y}_{n}$ satisfies the recursive identity for ${\bf z}_{n}=(\bm{y})$
 $Z\left(x_{n+1},{\bf z}_{n},\mu\right)$ $=$ $\sum_{a=1}^{g}\Theta_{a}(y_{n+1})\;O^{W_{\bm{\alpha}}}_{a}\left(z_{n+1};{\bf z}_{n}\right)$ (3.15)
@@
 Here $\partial^{(0,j)}$ is given by $\partial^{(i,j)}f(x,y)=\partial_{x}^{(i)}\partial_{y}^{(j)}f(x,y)$, for a function $f(x,y)$, and $\partial^{(0,j)}$ denotes partial derivatives with respect to $x$ and $y_{j}$.
-The forms $\Psi_{p}(y_{n+1},y_{k})\;dy_{k}^{j}$ given by (3.12), where $\Theta_{a}(x)$ is of (3.13), and $O^{W_{\bm{\alpha}}}_{a}(z_{n+1},{\bf z}_{n},\mu)$ of (3.14).
+The forms $\Psi_{p}(y_{n+1},y_{k})\;dy_{k}^{j}$ given by (LABEL:psih), where $\Theta_{a}(x)$ is of (LABEL:thetanew), and $O^{W_{\bm{\alpha}}}_{a}(z_{n+1},{\bf z}_{n},\mu)$ of (LABEL:oat).
 Acknowledgment